### v1

# Rethinking Gating Mechanism in Sparse MoE: Handling Arbitrary Modality Inputs with Confidence-Guided Gate

Abstract
Effectively managing missing modalities is a fundamental challenge in real-world multimodal learning scenarios, where data incompleteness often results from systematic collection errors or sensor failures. Sparse Mixture-of-Experts (SMoE) architectures have the potential to naturally handle multimodal data, with individual experts specializing in different modalities. However, existing SMoE approach often lacks proper ability to handle missing modality, leading to performance degradation and poor generalization in real-world applications. We propose Conf-SMoE to introduce a two-stage imputation module to handle the missing modality problem for the SMoE architecture and reveal the insight of expert collapse from theoretical analysis with strong empirical evidence. Inspired by our theoretical analysis, Conf-SMoE propose a novel expert gating mechanism by detaching the softmax routing score to task confidence score w.r.t ground truth. This naturally relieves expert collapse without introducing additional load balance loss function. We show that the insights of expert collapse aligns with other gating mechanism such as Gaussian and Laplacian gate. We also evaluate the proposed method on four different real world dataset with three different experiment settings to conduct comprehensive the analysis of Conf-SMoE on modality fusion and resistance to missing modality. We release our official implementation in: https://github.com/IcurasLW/Official-Repository-of-ConfSMoE.git
1 Introduction
As modern applications increasingly involve data from multiple heterogeneous sources, multimodal learning has emerged as a critical paradigm to unlock complementary insights and enhance decision-making in complex tasks spanning medical decision support[39, 41], cross-modal inference[31, 5], temporal reasoning[40, 23], and natural language understanding[9, 21]. Sparse Mixture-of-Experts (SMoE) architectures[34, 35, 22] are widely used in multimodal learning for their ability to assign different experts to different input patterns through conditional computation. They are particularly effective when all input modalities are present, enabling diverse feature extraction and flexible expert routing. However, this assumption is often violated in practice. Incomplete modality observations are common in real-world scenarios due to sensor failures, privacy restrictions, or heterogeneous data collection pipelines. When one or more modalities are missing, SMoE models receive partial input signals, which undermines the reliability of the gating mechanism and results in suboptimal expert selection. Although some methods attempt to assign different modality combinations to different experts, this strategy becomes computationally intractable as the number of modalities grows exponentially. To reduce this complexity, recent SMoE frameworks typically activate the Top-K experts per sample. However, the use of softmax-based routing[9] tends to produce sharp distributions, often concentrating the routing on a single expert. This leads to the expert collapse problem [9, 12, 35], where only a few experts dominate the computation as shown in Figure 1(a), resulting in poor diversity and limited generalization. Efforts to mitigate this issue using load-balancing losses introduce additional optimization challenges. In particular, they may cause unstable or ambiguous expert selection during training, as illustrated in Figure3(a).
(a) Expert Collapse
(b) Imputation by shared
(c) Ours
Figure 1: Expert Collapse of MoE and Comparison with Two Different Modality Imputation
While recent works[9, 12, 35] has tackled expert collapse within SMoE architectures, handling missing modalities remains an orthogonal and unresolved challenge. In the context of missing modalities, many recent approaches aim to reconstruct absent views through cross-modal information sharing [20, 29, 34] as shown in Figure 1(b). However, these methods typically introduce multiple auxiliary losses to jointly learn task-relevant, shared, and modality-specific features, which increases training complexity and may inadvertently erode modality-unique information. To alleviate this issue, recent work has proposed using learnable modality banks[35] to impute missing inputs based on the available modality combinations. Nevertheless, such banks are usually initialized randomly, similar to prompt tuning[15, 18], and often fail to capture fine-grained modality-specific structure, limiting their effectiveness in practical settings.
To overcome the limited robustness of existing SMoE models under incomplete modality conditions, we propose Confidence-guided Sparse Mixture-Of-E (Conf-SMoE), a unified framework that introduces two novel mechanisms for handling arbitrary missing modalities. Our method stems from a core insight: modality-specific structure should be preserved and leveraged for both reconstruction and routing, a property often overlooked in prior work. Based on this, we propose a two-stage imputation strategy that first estimates missing modalities using intra-modality correlations, and then refines these estimates through instance-level cross-modal alignment, as illustrated in Figure 1(c). The rationale behind this approach is inspired by the modality gap phenomenon discussed by [19], which suggests that a certain level of discrepancy between modalities is both expected and beneficial for downstream tasks. Thus, it is natural to first impute missing modalities using intra-modality information subsequently enhance them through instance-level cross-modal refinement.
We further uncover and formalize a critical cause of expert collapse in SMoE: the gradient concentration induced by softmax gating, which suppresses routing diversity and limits generalization[9, 30, 35, 12, 6]. Through theoretical analysis, we show that conventional load-balancing losses exacerbate this issue by introducing gradient conflicts during optimization, reflecting by the "Sinoual Wave" pattern in expert selection plot shown in Figure 3(a). To resolve this, we design a confidence-guided expert routing mechanism that replaces softmax gating with token-level confidence scores. This formulation not only avoids the adverse effects of entropy-based balancing but also enables stable, interpretable expert selection. By decoupling routing from the optimization bottlenecks of prior designs, Conf-SMoE achieves both enhanced expert diversity and consistent performance across a wide range of incomplete modality scenarios. Our contributions are the following:
  • A confidence-guided expert routing mechanism is proposed to mitigate expert collapse by decoupling gating from softmax-induced sharpness. Token-level confidence is used as a interpretable signal for expert selection.
  • A two-stage imputation framework is developed, where missing modalities are first reconstructed using modality-specific patterns, then refined with instance-level cross-modal context to preserve structural fidelity.
  • Empirical results across three missing-modality settings demonstrate strong robustness and generalization. Performance improvements are further supported by ablation studies that align with theoretical analysis.
2 Preliminaries
2.1 Mixture-of-Expert Architecture
In this work, we build upon the SMoE framework by incorporating expert routing into the Transformer’s feedforward layers, enabling dynamic token-to-expert assignment based on learned routing scores. This architecture supports selective expert activation, where only the Top-K experts are engaged per token, and proves especially effective in the joint expert setting[12], allowing tokens from different modalities to be routed to shared experts for enhanced cross-modal interaction. We begin with the standard Softmax SMoEs as a baseline configuration. Denote that input token embedding $\mathbf{h}\in\mathbb{R}^{d}$ and the a learnable softmax router $G(\mathbf{h})=softmax(\mathbf{h})$ that produces routing score $g_{i}$ for each expert $E_{i}$. Then, Top-K function is applied to select the top-k expert in forward, the final tokens/samples representation is weighted summation from Top-K expert with corresponding routing score $g_{i}$. Formally, we can represent the forward process of SMoE as:
$E(\mathbf{h})=\mathbf{h}+\sum_{i=1}^{K}G(\mathbf{h})E_{i}(\mathbf{h})$ (1)
Where $E_{i}(\cdot)$ represent the i-th expert chosen by the router in expert pool and we would like to sum up the representation from Top-K expert by gating score out of $G(\mathbf{h})$.
2.2 What Causes Expert Collapse?
The Jacobian matrix of Eq. 1 with respect to $\mathbf{h}$ can be analyzed to reveal the underlying cause of expert collapse in SMoE. Therefore,
$\mathbf{J}_{MoE}=\sum_{i=1}^{K}\frac{\partial softmax(\mathbf{h})}{\partial\mathbf{h}}E_{i}(\mathbf{h})+\sum_{i=1}^{K}G(\mathbf{h})E_{i}^{{}^{\prime}}(\mathbf{h})$ (2)
Let’s denote $\mathbf{g}=softmax(\mathbf{u})$, where $\mathbf{u}\in\mathbb{R}^{N}$ is the gating logits to select expert. The Jacobian matrix for the softmax function is denoted as $\mathbf{J}_{softmax}\in\mathbb{R}^{d\times d}$, where each element $\mathbf{J}_{ij}=\frac{\partial softmax(\mathbf{h_{j})}}{\partial\mathbf{h_{j}}}=g_{i}(\delta_{ij}-g_{i})$ and $\delta_{ij}$ is Kronecker delta. In matrix form, we can express Jacobian of softmax as: $\mathbf{J}_{softmax}=\textbf{diag(g)}-\textbf{g}\textbf{g}^{T}$, and the set of expert forward function $\mathbf{E}_{K}(\mathbf{h})=[E_{1}(\mathbf{h}),E_{2}(\mathbf{h}),\dots,E_{K}(\mathbf{h})]\in\mathbb{R}^{d\times d}$. For simplicity, we simplify the Top-K operation into $\mathbf{E_{K}}$. Therefore, we can rewrite Eq. 2 to Eq. 3: the first term is to learn a better routing score driven by gradient of gating score $\mathbf{g}$ and second term is to learn a better representation as the gradient driven by $E_{i}^{{}^{\prime}}(\mathbf{h})$. The Top-k mechanism makes the first term of Eq. 3 sparse and only a few experts contribute to gradient while only one expert dominate the gradient due to softmax sharp distribution. It is clear that the greatest gating score from the sharp distribution of softmax dominates the gradient in training phrase, resulting expert collapse based on token preference and produce rich-get-richer expert.
$\mathbf{J}_{MoE}=\underbrace{\mathbf{E}_{K}(\mathbf{h})(\textbf{ diag(g) }-\textbf{g}\textbf{g}^{\top})}_{\text{Learning a better routing score}}\;\;\;\;\;+\underbrace{\sum_{i=1}^{K}G(\mathbf{h})E_{i}^{{}^{\prime}}(\mathbf{h})}_{\text{Learning a better representation}}$ (3)
2.3 Gradient of Load Balance Loss
This expert collapse phenomenon restrict the model learning capacity since many other expert are wasted and cannot reflect the actual expert diversity. Two widely-used solutions to achieve balance expert usage is to introduce additional load balance loss and expert capacity. However, such load balance loss approach is difficult to optimize since it produces gradient direction opposite to softmax routing score terms in Eq 3. We consider that any load balance loss $\mathcal{L}_{load}$ will eventually results a flat and diverse softmax distribution, which encourage higher entropy on the softmax distribution. From Assumption 1 defined in Appendix, we define that the objective of any $\mathcal{L}_{load}$ as $\frac{1}{\mathcal{H(\mathbf{g})}}$ over the softmax distribution, where $\mathcal{H(\mathbf{g})}=-\sum_{i=1}^{N}g_{i}log\ g_{i}$ represent the entropy of softmax distribution g. Taking the Jacobian over the $\mathcal{L}_{load}$ will give us (See detailed derivation in Appendix. B.2):
$\mathbf{J}_{load}=[\frac{1}{\mathcal{H}(\textbf{g})^{2}}(log(\textbf{g})+1)^{\top}]\cdot(\textbf{diag(g)}-\textbf{g}\textbf{g}^{\top})$ (4)
Recall that $\mathbf{E}_{K}(\mathbf{h})$ in Eq. 3 is a set of expert with ReLU activation, we assume this network should not output negative value which is theoretically true in most cases. We know the fact that the Jacobian of softmax $\mathbf{J}_{softmax}$ is symmetry and always positive semi-definite[10], making the optimization of better routing score is always positive. In contrast, $log(\textbf{g})+1=[log(g_{1})+1,log(g_{2})+1,\dots,log(g_{K})+1]$. When sharp distribution appears, only one $g_{j}$ is closed to 1 while other $g_{i},(i\neq j)$ are closed to 0, which is practically and mostly true. Therefore, $\frac{1}{\mathcal{H(\textbf{g})}^{2}}(log(\textbf{g})+1)^{\top}$ is exponentially dominated by negative value and $\mathbf{J}_{load}$ is opposite to the first term of $\mathbf{J}_{MoE}$. Therefore, the opposite gradient from axillary load balance loss makes the training difficult to converge to global minimum and Expert selection becomes ambiguous, failing to reflect actual capacity of expert.
3 Propose Method
3.1 ConfNet:Confidence-Guided Gating Network
From Eq. 4 and Eq. 3, we understand that load balance loss shows opposite gradient direction to softmax router. Then, the potential solution to explicitly achieve load balance is to remove the association of sharp distribution from Eq. 3. One is to replace the routing function as illustrated in Figure. 2 (c) and one is to decouple the gating score with softmax router as illustrated in Figure. 2 (d). A typical example from a recent work, FuseMoE [12], they replace the Softmax router to Laplacian router and calculate the routing score based on similarity of the expert embedding $W_{N}$ by $\mathcal{L}_{1}$ distance. Then the Top-K experts are selected with highest similarity to expert embedding. The second approach is to decouple the connection of gating score and expert selection. One simple approach is to consider all expert are equally important, then the final representation is averaged from all selected expert. However, this averaged approach does not emphasize the difference of expert and lose the specification of expert.
Therefore, we propose a Confidence-Guided Gating Network Pool, which introduces a novel routing strategy guided by task confidence. Specifically, instead of solely relying on conventional softmax-based gating, which often results in sharp distributions and entangled gradient contributions, our method leverages the ground-truth label information to supervise the routing mechanism, encouraging expert selection aligned with task-relevant confidence rather than token or embedding similarity. For each expert $E_{i}$, we introduce an auxiliary confidence estimation network: $U_{i}$ that maps token embedding $\mathbf{h}\in\mathbb{R}^{d}$ to a single scalar logits $v_{i}\in\mathbb{R}$. $v_{i}$ is then mapped to a probability $g_{i}\in(0,1)$ as gating score by Sigmoid, representing the confidence of ground truth label $y_{t}$. To learn the true confidence $p_{t}$ of downstream task w.r.t $y_{t}$, where $t$ is the notation for ground truth. We apply MSE loss as an axillary loss that minimize the difference between $g_{i}$ and $p_{t}$. Formally, given any classification dataset $D$ with supervised signal $y_{i}$ and number of instance $|D|$, we consider the task loss $\mathcal{L}_{task}$ as Cross Entropy Loss and the final objective is to optimize:
$\mathcal{L}=\mathcal{L}_{task}+\mathcal{L}_{conf}=\frac{1}{|D|}\sum^{D}p_{t}log(p_{t})+\frac{1}{|D|K}\sum^{D}\sum^{K}_{i=1}(g_{i}-p_{t})^{2}$ (5)
The intuition is that the learning of gating score in Eq. 3 will be driven by the task-direct score, this forces model to select expert with task-preference instead of token preference as Softmax gating [9] or expert embedding preference as Laplacian gating does [12]. During inference, we use the learned confidence score $g_{i}$ from the ConfNet as a confidence proxy for the task-specific relevance of each expert. These scores are used for specialized token fusion, enabling dynamic integration of modality-specific and task-relevant information. The $g_{i}$ is Token-level confidence, which each token will have one corresponding $g_{i}$ for aggregation. Additionally, to capture coarse-grained signals, we aggregate token-level representations assigned to each expert and average them to compute expert-level gating signals as Expert-level variant, denoted as ConfSMoE-E in experiment. This confidence-driven routing enables both fine-grained specialization and robust multimodal interaction, leading to improved generalization and interpretability.
Figure 2: The Framework of Conf-SMoE: (a) Conf-SMoE consider the confidence learned from ConfNet pool as the gating score for expert aggregation. Missing modality imputation is achieved by taking the mean of random instances from the corresponding modality of training data as common modality specific feature. The instance-specific feature are sparsely aggregated from existing modality. (b) Expert collapse is because of the softmax sharp gating score. (c)(d) "Fair" gating score distribution and decoupling gating score can alleviate the expert collapse without the helps of load balance loss.
3.2 Missing Modality Modeling: Two-Stage Imputation
Inspired by the modality gap phenomenon observed in contrastive learning [19], we hypothesize that a similar gap exists in our multimodal setting—and that it may even contribute positively to model performance, as illustrated in Appendix Figure 4. The presence of this modality gap suggests that imputing a missing modality using only shared information from the remaining modalities is inherently challenging, due to the distinct representational spaces each modality occupies. Ideally, the imputed modality should preserve the characteristics of its original distribution without being significantly biased by other modalities. To address this, we propose a Two-stage Imputation strategy as following:
Pre-Imputation: When a modality $M_{i}\in\mathbb{R}^{d}$ is missing, we generate an pre-imputed modality embedding $\bar{M}_{i}$ by sampling $K$ instances from the modality pool $M_{i,n}=[M_{i,o},M_{i,1},\cdots,M_{i,n}]$, where $M_{i,n}$ represent the $i$-th modality from $n$-th instance modality pool. The modality pool contains all existing modality data for given number of modalities in training set. To remove the instance-specific feature and preserve the modality-common feature, we take an average over the $K$ instances as pre-imputed feature. The randomized selection of instances from the training set further aids in eliminating instance-specific noise during backpropagation. Thus, we send the existing and pre-imputed modality into backbone model to extract interaction between modality. Naturally, increasing $K$ yields more representative modality features but reduces stochasticity, potentially causing all imputed modalities to converge to the same representation across different missing-modality samples and the interaction between other modalities may be more deterministic and uninformative. We consider this $K$ as one of hyperparameter to finetune.
$\bar{M}_{i}=\frac{1}{K}\sum^{K}_{n=1}\mathbf{1}M_{i,n}$ (6)
Where $\mathbf{1}\in\{0,1\}^{n}$ is the selection vector for $M_{i}$ and $\mathbf{1}_{i}=1$ represents the $i$-th instance from modality pool is selected.
Post-Imputation: To add up instance-specific information from other modality, we propose a post-modality imputation to further capture the interactions after the MoE forward. Unlike conventional approaches that solely rely on the fused representation from one network, our method leverages information from multiple experts networks by selectively incorporating tokens beyond the Top-K experts. This allows the model to absorb nuanced, instance-specific features and interaction between modalities from different views of expert. To implement this, we employ a sparse attention mechanism, under the assumption that only a subset of modality-specific information is meaningfully shared across existing modalities. By enforcing sparsity, the model is encouraged to attend only to relevant parts of the available modalities, thereby reducing the risk of injecting modality-irrelevant or noisy information into the imputed representation, especially in heterogeneous or weakly correlated multimodal settings. Specifically, given a missing modality $M_{i}\in\mathbb{R}^{s\times d}$ and existing modality $M_{j}\in\mathbb{R}^{s\times d}$, where $s$ is the sequence length and $d$ is the hidden dimension. Each token of existing modality $t_{s}\in\mathbb{R}^{d}$ is sent to Top-K different experts to generate $K$ specialized tokens $t_{s,e}\in\mathbb{R}^{d}$, where $s$ represents the sequential position and $e$ represents expert index. We concatenate the specialized tokens $t_{s,e}$ in sequence dimension and apply cross attention operation over those specialized tokens and select only Top-K correlated specialized tokens for $\bar{M}_{i}$ to generate final imputed modality $M_{i}^{*}$.
$M_{i}^{*}=\bar{M}_{i}+SparseCrossAttention(\bar{M}_{i},M_{j},M_{j})$ (7)
We set the number of tokens $K$ selected from $t_{s,e}$ as $K=\frac{s(|M|-1)}{B}$, where $|M|$ represents the number of modality in dataset and $B$ is a sparsity hyperparameter, we set $B$ as 4 by default.
4 Experiment
4.1 Dataset
MIMIC-III [16] is a public dataset contains irregular time series and clinical notes modality. We follow the preprocessing scripts from MMIMI-III benchmark [13] and [38] to form multimodal patient instance. However, [38] filter the instances with missing modality and we modified the scripts to preserve with instances with missing modality. MIMIC-IV [17] is a public dataset contains irregular time series, clinical notes, ECG signal and chest image. We follow the preprocessing scripts from MedFuse [14] to preprocess irregular time series and further preprocess clinical notes, ECG signal and chest image by ourself. We follow the dataset split in MedFuse [14] and form multimodal patient instance. CMU-MOSI [36] is a multimodal sentiment analysis dataset comprising audio, video, and text modalities. CMU-MOSEI [2] is an extension of the CMU-MOSI dataset, utilizing the same modalities—audio, text, and video from YouTube recordings. We follow preprocessing script of CMU-MOSI and CMU-MOSEI from MMML[22].
4.2 Baselines
SMIL[20] impute missing modality with Bayesian meta-learning. ShaSpec[29] leverages multiple loss function to learn the modality-share and modality-specific information and impute the missing modality with modality-shared information. TF[37] extract the modality interaction by multimodal sub-networks and tensor fusion layer. mmFormer[28] is a typical multimodal transformer with multihead attention mechanism. LlMoE[22] achieves the multimodal learning with contrastive learning. FuseMoE[12] respect the missing modality and jointly learn single and missing modality combination data input and FuseMoE implement Laplacian gating mechanism for modality fusoin and expert selection. In experiment, we denote FuseMoE-S as softmax gating and FuseMoE-L as Laplacian gating. FlexMoE[35] assigns the combination of modality to different experts and impute the missing modality by modality bank. For our approach, we use ConfSMoE-T denotes Token-Level confidence fusion and ConfSMoE-E denotes Expert-level confidence fusion.
4.3 Experiment Setting
To ensure a fair comparison with baseline methods, we adopt the best-performing hyperparameter settings reported in their respective papers and official implementations. Additionally, we standardize the hidden dimension to 128 across all models. To eliminate confounding effects introduced by varying encoder architectures, we use the same modality-specific encoders for all models, as differences in encoder design can significantly influence final performance. For the MIMIC-III and MIMIC-IV datasets, we employ a pretrained ClinicalBERT[1] for textual data, a one-layer patch embedding module for both irregular time series and ECG signals, and a pretrained DenseNet[7] from TorchXRayVision for chest X-ray (CXR) images. Experiment Setting I: Natural Missing Modality. This setting reflects real-world missingness patterns observed in clinical datasets. Experiment Setting II: Random Modality Dropout. A fixed percentage of modalities are randomly dropped for each instance, while ensuring that each instance retains at least one modality in both train and test set. Experiment Setting III: Asymmetric Modality Dropout. Half of the modalities are randomly dropped during training, while testing is performed with only one or two modalities present per instance, simulating domain shifts or deployment constraints. All experiments are 3-Fold cross-validate in different seeds for fair comparison.
4.4 Primary Results
Experiment Setting I: In Table 1 and Appendix Table 9, we provide the experiments results on experiment setting I for three different ICU benchmarking task, where the modalities are natural missing instead of random. We obtain the following observations from the results: (1) In clinical benchmarking tasks 48-IHM, LOS, and 25-PHE, our proposed methods can significantly outperform the most recent baseline approaches such as FuseMoE and FlexMoE and other tradidtional multimodal learning methods in terms of F1 and AUC score. For MIMIC-III, the improvement on F1 and AUC is ranging from 1.81% to 3.91% and 0.92% to 1.22% respectively. (2) The improvement is even further when the size of dataset scale to bigger dataset but the same domain like MIMIC-IV. The improvement on F1 and AUC is ranging from 1.37% to 4.12% and 1.96% to 4.85%. The above empirical evidence roots and supports our claims that the model is strongly robust to missing modality and learn more comprehensive understanding on multimodal inputs.
Table 1: Experiment setting I: Main results on MIMIC-IV
Task Metric SMIL ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
48-IHM F1 39.58 $\pm$ 1.12 32.97 $\pm$ 1.08 45.39 $\pm$ 1.60 11.60 $\pm$ 0.54 43.76 $\pm$ 0.44 30.14 $\pm$ 1.15 40.21 $\pm$ 1.29 35.29 $\pm$ 0.30 49.18 $\pm$ 0.22 48.32 $\pm$ 0.30
AUC 76.60 $\pm$ 0.98 78.29 $\pm$ 0.25 80.43 $\pm$ 0.76 67.08 $\pm$ 0.77 82.97 $\pm$ 0.61 71.34 $\pm$ 0.88 78.05 $\pm$ 0.72 80.45 $\pm$ 0.44 85.24 $\pm$ 0.10 85.09 $\pm$ 0.17
LOS F1 58.52 $\pm$ 0.44 56.22 $\pm$ 0.28 57.06 $\pm$ 0.32 42.01 $\pm$ 1.12 59.03 $\pm$ 0.39 57.59 $\pm$ 0.76 58.31 $\pm$ 0.46 56.96 $\pm$ 0.53 61.33 $\pm$ 0.39 61.35 $\pm$ 0.41
AUC 75.86 $\pm$ 0.66 72.22 $\pm$ 0.18 74.65 $\pm$ 0.32 64.11 $\pm$ 0.40 76.17 $\pm$ 0.17 72.59 $\pm$ 1.07 72.59 $\pm$ 0.61 74.81 $\pm$ 0.37 78.22 $\pm$ 0.15 77.85 $\pm$ 0.12
25-PHE F1 27.77 $\pm$ 0.91 25.53 $\pm$ 0.11 26.15 $\pm$ 0.12 26.52 $\pm$ 0.29 25.38 $\pm$ 0.52 12.45 $\pm$ 0.54 12.25 $\pm$ 0.55 24.61 $\pm$ 0.71 28.67 $\pm$ 0.33 28.54 $\pm$ 0.25
AUC 63.90 $\pm$ 1.31 62.57 $\pm$ 0.23 70.33 $\pm$ 0.12 56.28 $\pm$ 0.16 72.50 $\pm$ 0.67 55.48 $\pm$ 0.39 58.52 $\pm$ 0.27 71.57 $\pm$ 0.47 74.56 $\pm$ 0.19 74.42 $\pm$ 0.34
Table 2: Experiment setting II: Main results on CMU-MOSI
Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
0% F1 41.78 $\pm$ 0.64 43.28 $\pm$ 0.36 50.22 $\pm$ 0.20 50.91 $\pm$ 0.19 47.69 $\pm$ 0.39 48.15 $\pm$ 0.57 51.50 $\pm$ 0.55 51.83 $\pm$ 0.27 51.70 $\pm$ 0.29
AUC 65.08 $\pm$ 0.39 64.48 $\pm$ 0.51 74.62 $\pm$ 0.49 78.12 $\pm$ 0.31 75.26 $\pm$ 0.20 75.96 $\pm$ 0.29 79.38 $\pm$ 0.23 80.56 $\pm$ 0.10 80.62 $\pm$ 0.07
10% F1 43.40 $\pm$ 0.74 40.14 $\pm$ 0.13 48.66 $\pm$ 1.20 49.91 $\pm$ 0.33 46.81 $\pm$ 0.45 48.81 $\pm$ 0.42 49.92 $\pm$ 0.28 49.99 $\pm$ 0.85 50.89 $\pm$ 0.07
AUC 65.45 $\pm$ 0.56 61.54 $\pm$ 0.66 77.21 $\pm$ 0.76 77.12 $\pm$ 0.14 70.25 $\pm$ 0.93 72.36 $\pm$ 0.50 77.07 $\pm$ 0.21 78.06 $\pm$ 0.64 78.03 $\pm$ 0.21
20% F1 39.15 $\pm$ 0.13 42.64 $\pm$ 0.30 48.51 $\pm$ 1.15 48.84 $\pm$ 0.61 47.92 $\pm$ 0.58 45.00 $\pm$ 0.44 49.16 $\pm$ 0.24 50.58 $\pm$ 0.73 50.17 $\pm$ 0.31
AUC 60.87 $\pm$ 0.29 65.52 $\pm$ 0.98 74.16 $\pm$ 0.71 76.64 $\pm$ 0.23 71.05 $\pm$ 0.36 72.15 $\pm$ 0.37 73.80 $\pm$ 1.43 77.17 $\pm$ 0.55 77.18 $\pm$ 0.11
30% F1 39.54 $\pm$ 0.17 40.42 $\pm$ 0.75 46.66 $\pm$ 1.24 46.13 $\pm$ 0.55 43.78 $\pm$ 0.51 44.48 $\pm$ 0.63 45.49 $\pm$ 0.44 47.57 $\pm$ 0.10 47.52 $\pm$ 0.77
AUC 60.80 $\pm$ 0.44 60.57 $\pm$ 0.99 71.25 $\pm$ 0.57 74.33 $\pm$ 0.49 70.38 $\pm$ 0.56 72.25 $\pm$ 0.48 72.54 $\pm$ 0.92 75.59 $\pm$ 0.18 75.42 $\pm$ 0.49
40% F1 35.32 $\pm$ 0.32 38.69 $\pm$ 0.42 42.22 $\pm$ 0.44 43.31 $\pm$ 0.38 41.88 $\pm$ 0.42 43.20 $\pm$ 0.20 43.13 $\pm$ 0.61 45.12 $\pm$ 0.27 44.18 $\pm$ 0.10
AUC 59.58 $\pm$ 0.13 61.62 $\pm$ 0.07 68.44 $\pm$ 0.14 67.21 $\pm$ 0.43 64.61 $\pm$ 0.63 66.14 $\pm$ 0.11 68.12 $\pm$ 0.34 71.65 $\pm$ 0.47 69.55 $\pm$ 0.49
50% F1 36.39 $\pm$ 0.97 38.46 $\pm$ 1.09 41.14 $\pm$ 0.31 41.54 $\pm$ 0.77 38.55 $\pm$ 0.34 40.90 $\pm$ 0.49 41.95 $\pm$ 0.48 43.89 $\pm$ 0.51 42.48 $\pm$ 0.12
AUC 55.60 $\pm$ 0.12 60.49 $\pm$ 0.80 61.87 $\pm$ 0.67 67.43 $\pm$ 0.34 62.28 $\pm$ 0.28 65.73 $\pm$ 0.30 66.47 $\pm$ 0.49 71.51 $\pm$ 0.53 69.04 $\pm$ 0.32
Experiment Setting II: In Table 2 and Appendix E Table 10, we implement Random Modality Dropout to study the sensitivity of model to missing modality. In this setting, we empirically find that those methods with missing modality imputation usually shows higher tenacity in high proportion of missingness such as FlexMoE and our propose method. Although ShaSpec has also impute the missing modality by shared information, but it is weak on the backbone model and struggling to optimize with multi-task objectives. Our proposed method, on the other hand, shows superior performance across all setting in CMU-MOSEI and CMU-MOSI for all metrics.
Experiment Setting III: In Appendix E Table 11, Table 12, Table 13 and Table 14, we study which modality combination is significant and how is the tenacity of our proposed method to different distribution shift on missing modality. We find that single modality shows very similar performance during inference phrase in our proposed method and Flex-MoE across AUC and F1, indicating the effectiveness of missing modality imputation. In particular, our proposed also shows the highest resistance to missing modality across different combination and remains competitive performance no matter what modality is missing. In addition, the modality combination Audio-Video is the worse modality combination for all implemented models. The potential reason may be that the data interaction between those two modalities are weak or even opposite. We show some visualization evidence in Figure 4 that the embedding of those two modalities are mostly symmetry.
Table 3: Ablation I: Module dropout. CMU-MOSI adopt 50% missing in Experiment III
MIMIC-III CMU-MOSI
F1 AUC F1 AUC
w/o impute 43.79 $\pm$ 2.81 85.33 $\pm$ 1.22 42.51 $\pm$ 2.25 69.67 $\pm$ 1.27
w/o Conf 48.59 $\pm$ 1.97 85.91 $\pm$ 0.84 42.47 $\pm$ 0.94 68.77 $\pm$ 1.66
w/o impute & Conf 46.32 $\pm$ 1.66 85.15 $\pm$ 0.57 42.50 $\pm$ 0.63 66.73 $\pm$ 1.41
Conf-SMoE 53.44 $\pm$ 0.27 87.05 $\pm$ 0.58 43.89 $\pm$ 0.51 71.51 $\pm$ 0.53
4.5 Are Missing Modality Imputation and ConfNet Actually Useful?
We further implement Ablation Study I in Table 3 by dropping off the module we designed in proposed method and study which part of of model is actually useful. Specifically, we dropout the modality imputation module and keep the missing modality as zeros, we can see that F1 and AUC dropped off by 6.21% and 1.81% on MIMIC-III, 5.62% and 0.81% on CMU-MOSI. This performance drops is even worse when imputation and ConfNet gating mechanism is removed from proposed method, providing strong empirical evidence for effectiveness of our design. It is interesting that the w/o impute & Conf variant outperform w/o impute, indicating the model may only focus on existing modality instead of modality interaction.
4.6 Why Does Confidence-Guided Gating Improve Performance?
We implemented experiment on switching different gating mechanisms in our propose method as shown in Table 4 and we also plot the expert selection heatmap in Appendix F. Softmax gating mechanism has the worse expert collapse, leading to the poorer performance than the balance softmax gating. However, such load balance loss is hard to learn proper routing scores since the optimization is opposite as stated in Section 2, leading to suboptimal solution in this ablation. Softmax w/ $\mathcal{L}_{load}$ variant shows "Sharp Sinusoidal Wave" expert selection pattern as shown in Figure 3(a), indicating the selection is ambiguous and difficulty on convergence. We also found that considering the uniform weight for all experts can achieve similar performance as sole softmax and generate similar sinusoidal pattern as softmax w/ $\mathcal{L}_{load}$, shown in Figure 9.
(a) Expert selection of softmax w/ $\mathcal{L}_{laod}$
(b) Expert selection of ConfNet
Figure 3: Expert selection: X-axis represents epoch and Y-axis represents expert ID. The darker the color, the more selection will be for a particular expert. See detailed and more plots in Appendix F
Table 4: Ablation II: Different router. CMU-MOSI adopt 50% missing setting in Experiment III
Task Metric Mean Softmax Softmax w/ $\mathcal{L}_{load}$ Gaussian Laplacian ConfNet
MIMIC-III F1 48.34 $\pm$ 0.22 48.59 $\pm$ 1.97 51.67 $\pm$ 0.62 46.50 $\pm$ 1.62 50.10 $\pm$ 1.17 53.44 $\pm$ 0.27
AUC 85.13 $\pm$ 0.16 85.91 $\pm$ 0.84 85.97 $\pm$ 0.52 86.70 $\pm$ 0.33 85.47 $\pm$ 1.05 87.05 $\pm$ 0.58
CMU-MOSI F1 42.29 $\pm$ 0.87 41.47 $\pm$ 0.94 42.43 $\pm$ 1.34 42.57 $\pm$ 1.81 43.15 $\pm$ 1.53 43.89 $\pm$ 0.51
AUC 68.01 $\pm$ 1.68 68.77 $\pm$ 1.66 67.40 $\pm$ 0.54 67.74 $\pm$ 0.45 67.60 $\pm$ 1.54 71.51 $\pm$ 0.53
In addition, Laplacian and Gaussian gating function can achieve comparable to softmax w/ $\mathcal{L}_{load}$ gating mechanism without helps of $\mathcal{L}_{load}$, providing further evidence to support the advantage of balance expert selection. Our proposed ConfNet without helps of $\mathcal{L}_{load}$ is the best and significantly outperform other gating function in CMU-MOSI and MIMIC-III. As illustrated in Figure 6 and supported by the quantitative results in Table 4, ConfNet not only preserves expert specialization but also achieves better load balancing than Laplacian, albeit slightly less balanced than softmax with $\mathcal{L}_{\text{load}}$ and Gaussian gating as shown in Appendix F. This comparison provides two insights: First, optimization of $\mathcal{L}_{load}$ is simply against to optimization of softmax gating score but does not help to identify the best expert, reflecting the "Sharp Sinusoidal Wave" pattern in expert selection and ambiguous expert selection. Second, learning specialization of expert can also benefit to downstream performance. Notably, both ConfNet and Laplacian gates demonstrate a common pattern: they maintain the specialization of dominant experts while redistributing some load to less active ones, which highlights that strict load balancing may suppress expert specialization and degrade performance, whereas maintaining a balance between specialization and load distribution leads to superior results.
4.7 Complexity Analysis
Table 5: Complexity analysis on CMU-MOSI with 50% missing setting in Experiment III
Metrics ShaSpec TF LlMoE mmFormer FuseMoE FlexMoE ConfSMoE
MFLOPs 0.51 8.42 1016.15 1577.63 199.80 20.16 47.10
#Params 102,022 4,211,331 3,457,819 11,373,315 243,543,180 1,126,147 3,116,310
F1 score 36.39 $\pm$ 0.97 38.46 $\pm$ 1.09 41.54 $\pm$ 0.77 41.14 $\pm$ 0.31 40.90 $\pm$ 0.49 41.95 $\pm$ 0.30 43.89 $\pm$ 0.51
In Table 5, we compared the complexity between our proposed method and baseline models in terms of FLOPs and number of parameters. ShaSpec is the simplest models, but it turns out not comparable performance in many tasks. On the other hand, those MoE-based approach like FuseMoE and LlMoE are extremely complicated either in FLOPs and number of parameters while the performance is suboptimal compare to FlexMoE and our proposed method. mmFormer show very comparable results with our proposed method, however, the FLOPs and parameters are extremely larger than ConfSMoE, leading to poor training and inference efficiency. On the other hand, FlexMoE is 2 times less complicated than ConfSMoE while the performance is worse than LlMoE, mmFormer and ConfSMoE. By balancing complexity and performance from Table 5, it is clear that ConfSMoE is the best model to balance complexity and performance.
5 Related Works
While in many works considers that the interaction between multimodal data can bring comprehensive understanding for downstream task. Initial multimodal fusion approach usually incoperates neural kernel fusion[4, 24], early/late network fusion [32, 3, 11]. Approaches such as Tensor Fusion Network[37], Multimodal Transformer[28] and Multimodal Adaption Gate[25], LlMoE[22] all highlights the multimodal interactions are beneficial to downstream task. In medical domain, MedFuse [14] incorporated CXR and EHR modality to jointly fuse modality with early and later fusion in a single fusion network. MISTS [38] incorporated multi-time attention mechanism to encoder irregular time serires and adapts multimodal cross attention to fuse two different modality. [33] fused multimodal data in a attention gate and compute a replacement vector for modality fusion. However, these approaches are limited to full modality only and difficult to extend to increasing number of modality since the complexity can be significantly increased by the number of modality. Missing modality is realistic and practical problem in real life application such as medical diagnose. SMIL [20] proposed a Bayesian meta-learning solution with reconstruction network to impute the missing modality from in hidden state during fusion. [8, 26] propose generative adversarial networks to impute the missing modality by semi-supervised learning. ShaSpec [29] and DrFuse [34] designed multiple loss function to learn share and specific representation of different modality. However, the reconstruction network from SMIL and ShaSpec requires multiple additional auxiliary loss function, which occupies the great amount of gradient in optimization process instead of task specific loss. [27] In addition, recent studies like FuseMoE [12] respect the missingness of modalities and wish to assign each combination of multimodal data to different experts. Flex-MoE[35] further expand the advantages of FuseMoE and introduce a modality bank to impute the missing modality.
6 Conclusion
In this work, a novel Conf-SMoE model, a confidence-guided sparse mixture-of-experts framework, is proposed to robustly handle multimodal learning with missing modalities. By theoretically analyzing the root cause of expert collapse and empirically validating our findings, we propose two key innovations: a two-stage imputation strategy that preserves modality-specific structure, and a novel confidence-driven gating mechanism that decouples expert selection from softmax-induced sharpness. These components jointly improve both expert specialization and routing stability without requiring entropy-based balancing losses.
Appendix A Notation
To facilitate understanding of the model formulation and derivations, Table 6 outlines the notations used throughout this paper. These include variables associated with expert routing, gating mechanisms, modality assignments, and sparsity control, which are central to our proposed Conf-SMoE framework.
Table 6: Notation
Variable Definition
$E_{i}$ i-th Expert Forward Function
$\mathbf{h}$ Embedding of Arbitrary Data Input
$K$ Number of K selection
$N$ Number of Expert
$G(\cdot)$ Gating Function
$\mathbf{g}$ Gating Score Vector
$\mathcal{H}(\cdot)$ Information Entropy
$p_{t}$ Confidence of Ground Truth
$M_{i,n}$ Modality Pool with i instance and n modality
$t_{s,e}$ Tokens at sequential position $s$ and assigned expert with index $e$
$B$ Sparsity Hyperparameter
$|M|$ Number of Modality
$\mathbf{diag(\cdot)}$ Diagonal Matrix
$\delta_{ij}$ Kronecker delta
$u_{i}$ softmax input at entry i
Appendix B Derivation of Jacobian for MoE and Load Balance Loss
B.1 Jacobian for Softmax
This section provides detailed mathematical derivations of the Jacobian matrix used in the gating function of Conf-SMoE and its relation to the load balance loss. We first derive the Jacobian of the softmax operation applied to the gating logits, which is essential for analyzing expert selection dynamics. We then extend this analysis to derive gradients of the entropy-based load balance loss. These derivations provide theoretical insights into how sparse expert routing and balanced expert utilization can be jointly optimized.
We first derive the Jacobian of the softmax function as used in the MoE formulation Eq. 1.
$\textbf{g}=softmax(u)=\frac{e^{u_{i}}}{\sum^{N}_{j=1}e^{u_{j}}}$ (8)
where $u_{i}$ is the i-th input value of gate logits.
$\frac{\partial\textbf{g}}{\partial u}=\left\{\begin{aligned} \frac{e^{u_{i}}(\sum^{N}_{j=1}e^{u_{j}}-e^{u_{i}})}{(\sum^{N}_{j=1}e^{u_{j}})^{2}}&\quad\text{if }i=j\\ -\frac{e^{u_{i}}e^{u_{j}}}{(\sum^{N}_{j=1}e^{u_{j}})^{2}}&\quad\text{if }i\neq j\end{aligned}\right.$ (9)
$\frac{\partial\textbf{g}}{\partial u}=\left\{\begin{aligned} g_{i}(1-g_{j})&\quad\text{if }i=j\\ -g_{i}g_{j}&\quad\text{if }i\neq j\end{aligned}\right.$ (10)
Where $g_{i}$ represent $\frac{e^{u_{i}}}{\sum^{N}_{j=1}e^{u_{j}}}$. This can be further compactly represent as:
$\frac{\partial\textbf{g}}{\partial u}=g_{i}(\delta_{ij}-g_{j})$ (11)
Where $\delta_{ij}=\begin{cases}1&\text{if }i=j\\ 0&\text{if }i\neq j\end{cases}$ is Kronecker delta and the matrix form is:
$\textbf{J}_{softmax}=\frac{\partial\textbf{g}}{\partial u}=\textbf{diag(g)}-\textbf{g}\textbf{g}^{\top}$ (12)
B.2 Jacobian for Load Balance Loss
Assumption 1
For any load balance loss $\mathcal{L}_{\text{load}}$ applied to a Softmax-based MoE, the effect of $\mathcal{L}_{\text{load}}$ is to encourage a flat expert selection distribution, which is equivalent to maximizing the entropy of the gating scores.
With Assumption 1, We have load balance loss $\mathcal{L}_{load}=\frac{1}{\mathcal{H}(\textbf{g})}$, where $\mathcal{H}(\textbf{g})=-\sum^{N}_{i=1}g_{i}log(g_{i})$ is the entropy of the gating distribution.
By applying the chain rule, we obtain:
$\frac{\partial\mathcal{L}_{load}}{\partial\textbf{u}}=-\frac{1}{\mathcal{H}(\textbf{u})^{2}}\frac{\partial\mathcal{H}}{\partial\textbf{u}}=-\frac{1}{\mathcal{H}(\textbf{u})^{2}}\frac{\partial\mathcal{H}}{\partial\textbf{g}}\frac{\partial\textbf{g}}{\partial\textbf{u}}$ (13)
The intermediate gradients are given by:
$\frac{\partial\mathcal{H}}{\partial\textbf{g}}=-log(\textbf{g})-\textbf{g}\frac{1}{\textbf{g}}=-log(\textbf{g})-1,\quad\frac{\partial\textbf{g}}{\partial\textbf{u}}=\textbf{diag(g)}-\textbf{g}\textbf{g}^{\top}$ (14)
Therefore, he final Jacobian of the load balance loss $J_{load}$ is:
$\mathbf{J}_{load}=\frac{\partial\mathcal{L}_{load}}{\partial\textbf{g}}=[\frac{1}{\mathcal{H}(\textbf{g})^{2}}(log\textbf{g}+1)^{\top}]\cdot(\textbf{diag(g)}-\textbf{g}\textbf{g}^{\top})$ (15)
Appendix C Experiment Details
C.1 Dataset and Task Description
Table 7: Dataset statistics
Dataset Task #Samples Modality #Classes Is Imbalanced? Missing Ratio
MIMIC-III 48-IHM 6621 Time Series, Text 2 Yes 24.21%
LOS 6621 Time Series, Text 2 Yes 24.21%
25-PHE 12,278 Time Series, Text 2 Yes 24.25%
MIMIC-IV 48-IHM 22,733 Time Series, Text 2 Yes 47.10%
LOS 22,733 Time Series, Text 2 Yes 47.10%
25-PHE 47,027 Time Series, Text 2 Yes 50.32%
CMU-MOSI Sentiment Analysis 1870 Audio, Video, Text 3 Yes 0% - 50%
CMU-MOSEI Sentiment Analysis 20,985 Audio, Video, Text 3 Yes 0% - 50%
48-IHM is a binary classification task, where the time series is truncated to 48 hours readings and predict whether the patient in-hospital mortality in ICU.
LOS is also binary classification task. We formulate task similar to 47-IHM and those patients who spent at least 48 hours in the ICU are filtered to predict the remaining Length Of Stay.
25-PHE is a multilabel classification problem. We attempt to predict one 25 acute care conditions presented in a givem ICU stay record. The definition of those 25 acute care conditions are defined in MIMIC benchmark [16]
CMU-MOSI and CMU-MOSEI are multimodal sentiment analysis benchmarks consisting of video clips annotated with sentiment scores. Each sample includes aligned audio, video, and text modalities. To simulate modality missingness, we introduce controlled random dropout during training.
All datasets are publicly available and widely adopted in multimodal research. Those datasets can be found and downloaded from the following links:
  • CMU-MOSI and CMU-MOSEI download link: https://github.com/zehuiwu/MMML
  • MIMIC-III download link: https://physionet.org/content/mimiciii/1.4/ and benchmark: https://github.com/YerevaNN/mimic3-benchmarks
  • MIMIC-IV download link: https://physionet.org/content/mimiciv/2.2/ and benchmark: https://github.com/nyuad-cai/MedFuse
C.2 Hyperparameter and Hardware Setting
All experiments are conducted in a remote server with 4 $\times$ RTX4090 24GB with 256 RAM and 96-cores CPU.
Table 8: Hyperparameter setting
Parameter Value
#Expert 8 for MIMIC-III and MIMIC-IV, 4 for CMU-MOSI and CMU-MOSEI
Top-K Expert 2
Learning Rate 3e-4
Hidden Size 128
#MoE layers 1
Training Epoch 50
Weight for $\mathcal{L}_{conf}$ 1
# K for Pre-imputation 10
Random Seed 2023, 2024, 2025
Dropout 0.1
Appendix D Limitation
This work may be limited to the task with supervision signal and additional RAM memory usage is required since the modality pool requires additional space for saving modality data.
Appendix E Additional Results
These experiments extend the main results presented in Section Experiment and highlight the scalability of our method to real-world multimodal challenges.
Table 9: Experiment setting I: Main results on MIMIC-III
Task Metric SMIL ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
48-IHM F1 48.56 $\pm$ 0.35 34.36 $\pm$ 0.73 43.80 $\pm$ 0.67 23.09 $\pm$ 1.13 47.62 $\pm$ 0.21 48.62 $\pm$ 0.27 51.63 $\pm$ 0.39 47.81 $\pm$ 0.31 53.44 $\pm$ 0.27 50.00 $\pm$ 2.17
AUC 84.00 $\pm$ 1.10 79.86 $\pm$ 0.53 81.27 $\pm$ 0.76 78.47 $\pm$ 0.65 83.30 $\pm$ 0.64 85.81 $\pm$ 0.13 86.22 $\pm$ 0.15 85.10 $\pm$ 0.61 87.05 $\pm$ 0.58 87.14 $\pm$ 0.21
LOS F1 62.01 $\pm$ 0.54 57.27 $\pm$ 0.36 61.10 $\pm$ 0.65 56.23 $\pm$ 0.79 62.75 $\pm$ 0.10 62.26 $\pm$ 0.36 62.27 $\pm$ 1.61 64.18 $\pm$ 0.34 66.27 $\pm$ 0.40 66.39 $\pm$ 0.23
AUC 80.08 $\pm$ 0.35 74.71 $\pm$ 0.07 77.94 $\pm$ 0.12 72.60 $\pm$ 0.79 79.80 $\pm$ 0.27 80.05 $\pm$ 0.51 79.70 $\pm$ 0.43 80.49 $\pm$ 0.48 81.66$\pm$ 0.26 81.71 $\pm$ 0.18
25-PHE F1 33.01 $\pm$ 0.30 19.52 $\pm$ 0.39 35.10 $\pm$ 0.38 30.38 $\pm$ 0.38 36.22 $\pm$ 0.39 33.48 $\pm$ 0.22 34.07 $\pm$ 0.51 35.31 $\pm$ 0.57 37.40 $\pm$ 0.13 40.13 $\pm$ 0.40
AUC 75.17 $\pm$ 1.84 64.61 $\pm$ 0.25 71.79 $\pm$ 0.26 66.56 $\pm$ 0.17 77.93 $\pm$ 0.44 77.22 $\pm$ 0.29 77.19 $\pm$ 0.33 77.37 $\pm$ 0.10 78.93 $\pm$ 0.38 78.28 $\pm$ 0.73
Table 10: Experiment setting II: Main results on CMU-MOSEI
Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
0% F1 47.48 $\pm$ 0.27 38.89 $\pm$ 0.54 56.75 $\pm$ 0.47 59.13 $\pm$ 0.52 57.04 $\pm$ 0.15 58.71 $\pm$ 0.23 60.62 $\pm$ 0.96 61.31 $\pm$ 0.14 62.35 $\pm$ 0.12
AUC 67.93 $\pm$ 0.56 64.26 $\pm$ 0.61 76.42 $\pm$ 0.24 80.29 $\pm$ 0.31 78.54 $\pm$ 0.42 78.13 $\pm$ 0.34 81.34 $\pm$ 0.20 82.87 $\pm$ 0.10 82.75 $\pm$ 0.44
10% F1 47.41 $\pm$ 0.18 34.94 $\pm$ 1.10 55.37 $\pm$ 0.47 57.46 $\pm$ 1.10 55.90 $\pm$ 0.55 56.68 $\pm$ 0.31 58.66 $\pm$ 1.10 59.81 $\pm$ 0.34 60.22 $\pm$ 0.31
AUC 69.26 $\pm$ 0.22 61.05 $\pm$ 0.34 75.36 $\pm$ 0.53 79.38 $\pm$ 0.94 76.38 $\pm$ 0.39 76.44 $\pm$ 0.12 79.48 $\pm$ 0.23 81.00 $\pm$ 0.10 81.05 $\pm$ 0.17
20% F1 46.54 $\pm$ 1.04 35.37 $\pm$ 0.94 52.91 $\pm$ 0.75 55.35 $\pm$ 0.88 54.44 $\pm$ 0.27 54.60 $\pm$ 0.19 55.96 $\pm$ 1.67 58.06 $\pm$ 0.21 57.66 $\pm$ 0.20
AUC 68.05 $\pm$ 0.56 59.54 $\pm$ 0.22 73.02 $\pm$ 0.65 77.29 $\pm$ 0.47 74.32 $\pm$ 0.16 74.92 $\pm$ 0.64 78.05 $\pm$ 0.10 78.85 $\pm$ 0.28 78.83 $\pm$ 0.22
30% F1 45.05 $\pm$ 0.90 35.36 $\pm$ 0.84 49.71 $\pm$ 0.30 53.93 $\pm$ 1.13 52.27 $\pm$ 0.35 49.92 $\pm$ 0.36 52.00 $\pm$ 0.47 54.16 $\pm$ 0.49 54.83 $\pm$ 0.15
AUC 66.89 $\pm$ 0.31 60.23 $\pm$ 0.31 71.28 $\pm$ 0.10 76.50 $\pm$ 0.31 72.22 $\pm$ 0.52 72.52 $\pm$ 0.29 76.14 $\pm$ 0.38 76.91 $\pm$ 0.41 77.10 $\pm$ 0.30
40% F1 43.49 $\pm$ 0.95 32.23 $\pm$ 0.61 48.06 $\pm$ 0.54 50.95 $\pm$ 0.99 49.32 $\pm$ 0.17 49.36 $\pm$ 0.60 49.29 $\pm$ 0.28 50.72 $\pm$ 0.37 50.66 $\pm$ 0.18
AUC 65.32 $\pm$ 0.59 60.09 $\pm$ 0.35 69.31 $\pm$ 0.10 73.24 $\pm$ 0.76 69.85 $\pm$ 0.11 70.05 $\pm$ 0.46 72.08 $\pm$ 1.27 73.63 $\pm$ 0.48 73.88 $\pm$ 0.25
50% F1 42.75 $\pm$ 1.22 33.01 $\pm$ 0.51 47.39 $\pm$ 0.16 46.69 $\pm$ 0.63 43.08 $\pm$ 0.47 $4$$\pm$ 0.36 46.19 $\pm$ 0.34 47.31 $\pm$ 0.20 48.13 $\pm$ 0.33
AUC 62.32 $\pm$ 0.42 58.20 $\pm$ 0.17 66.61 $\pm$ 0.22 70.90 $\pm$ 0.82 67.63 $\pm$ 0.55 67.44 $\pm$ 0.16 70.15 $\pm$ 1.10 71.74 $\pm$ 0.14 70.48 $\pm$ 0.11
Table 11: Experiment setting III: Main results on CMU-MOSI
Test Modality Dataset: CMU-MOSI / Evaluation Metric: AUC
Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
62.18 $\pm$ 0.88 72.72 $\pm$ 0.08 59.37 $\pm$ 1.12 74.63 $\pm$ 0.13 55.35 $\pm$ 1.11 56.44 $\pm$ 1.48 74.59 $\pm$ 1.50 75.57 $\pm$ 0.55 75.56 $\pm$ 0.33
55.79 $\pm$ 2.53 52.21 $\pm$ 0.31 57.87 $\pm$ 0.32 56.84 $\pm$ 0.34 56.24 $\pm$ 1.96 55.11 $\pm$ 0.77 52.28 $\pm$ 1.23 58.42 $\pm$ 0.14 58.66 $\pm$ 0.20
65.22 $\pm$ 1.01 71.96 $\pm$ 0.47 57.89 $\pm$ 0.65 75.59 $\pm$ 0.39 54.36 $\pm$ 1.67 56.10 $\pm$ 1.76 75.55 $\pm$ 1.64 77.02 $\pm$ 0.63 76.96 $\pm$ 0.64
55.10 $\pm$ 1.59 71.01 $\pm$ 0.53 58.57 $\pm$ 0.44 74.73 $\pm$ 0.51 55.94 $\pm$ 1.51 55.66 $\pm$ 1.66 74.35 $\pm$ 1.60 75.07 $\pm$ 0.70 75.11 $\pm$ 0.54
61.60 $\pm$ 4.80 73.77 $\pm$ 0.44 60.28 $\pm$ 1.42 75.73 $\pm$ 0.33 55.45 $\pm$ 0.95 56.58 $\pm$ 1.29 74.42 $\pm$ 1.39 75.58 $\pm$ 0.55 75.50 $\pm$ 0.41
66.85 $\pm$ 3.01 69.04 $\pm$ 0.94 57.19 $\pm$ 0.72 73.29 $\pm$ 0.53 56.04 $\pm$ 1.91 57.85 $\pm$ 0.74 75.34 $\pm$ 1.46 77.02 $\pm$ 0.63 76.47 $\pm$ 0.63
Table 12: Experiment setting III: Main results on CMU-MOSI
Test Modality Dataset: CMU-MOSI / Evaluation Metric: F1
Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
41.27 $\pm$ 0.81 48.25 $\pm$ 0.11 34.96 $\pm$ 0.72 44.45 $\pm$ 0.33 30.51 $\pm$ 2.53 29.90 $\pm$ 2.58 50.54 $\pm$ 0.73 54.07 $\pm$ 0.29 54.04 $\pm$ 0.36
36.99 $\pm$ 2.03 35.79 $\pm$ 0.55 34.83 $\pm$ 0.54 35.20 $\pm$ 0.64 24.29 $\pm$ 4.23 27.89 $\pm$ 6.23 30.02 $\pm$ 7.82 37.40 $\pm$ 0.16 38.17 $\pm$ 0.13
43.72 $\pm$ 1.64 45.53 $\pm$ 0.33 37.25 $\pm$ 0.53 47.79 $\pm$ 0.46 20.68 $\pm$ 2.14 29.88 $\pm$ 4.55 50.41 $\pm$ 1.55 52.86 $\pm$ 0.75 51.13 $\pm$ 0.15
42.98 $\pm$ 0.21 46.49 $\pm$ 0.91 38.11 $\pm$ 0.35 44.86 $\pm$ 0.94 25.67 $\pm$ 2.82 27.34 $\pm$ 2.94 50.39 $\pm$ 0.87 53.76 $\pm$ 0.43 54.57 $\pm$ 0.35
41.36 $\pm$ 2.27 47.13 $\pm$ 0.54 37.99 $\pm$ 0.40 43.81 $\pm$ 0.42 25.70 $\pm$ 2.82 27.30 $\pm$ 2.95 50.65 $\pm$ 0.90 53.77 $\pm$ 0.45 49.93 $\pm$ 0.06
45.45 $\pm$ 2.73 46.02 $\pm$ 1.21 38.79 $\pm$ 0.56 45.50 $\pm$ 0.47 25.23 $\pm$ 4.59 29.73 $\pm$ 3.51 50.57 $\pm$ 1.35 53.72 $\pm$ 0.94 50.69 $\pm$ 0.47
Table 13: Experiment setting III: Main results on CMU-MOSEI
Test Modality Dataset: CMU-MOSEI / Evaluation Metric: AUC
Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
64.47 $\pm$ 0.10 76.60 $\pm$ 0.39 56.98 $\pm$ 0.20 81.23 $\pm$ 0.07 75.81 $\pm$ 0.82 76.62 $\pm$ 0.37 80.63 $\pm$ 0.41 81.45 $\pm$ 0.02 81.52 $\pm$ 0.07
58.41 $\pm$ 0.10 56.72 $\pm$ 0.84 58.41 $\pm$ 0.17 59.64 $\pm$ 0.56 57.32 $\pm$ 0.62 57.22 $\pm$ 0.12 58.28 $\pm$ 1.51 60.33 $\pm$ 0.03 60.00 $\pm$ 0.23
64.46 $\pm$ 0.14 75.91 $\pm$ 0.04 58.64 $\pm$ 0.13 80.00 $\pm$ 0.29 77.24 $\pm$ 0.60 77.31 $\pm$ 0.16 80.00 $\pm$ 0.10 81.26 $\pm$ 0.10 81.23 $\pm$ 0.14
63.52 $\pm$ 0.19 75.82 $\pm$ 0.28 58.71 $\pm$ 0.35 80.69 $\pm$ 0.34 76.84 $\pm$ 0.14 76.27 $\pm$ 0.70 81.06 $\pm$ 0.20 81.60 $\pm$ 0.13 81.53 $\pm$ 0.22
64.40 $\pm$ 0.24 75.83 $\pm$ 0.28 58.89 $\pm$ 0.40 80.96 $\pm$ 0.26 76.85 $\pm$ 0.14 76.51 $\pm$ 0.50 80.48 $\pm$ 0.21 81.49 $\pm$ 0.13 81.65 $\pm$ 0.17
65.16 $\pm$ 0.04 75.85 $\pm$ 0.09 58.32 $\pm$ 0.04 80.19 $\pm$ 0.11 77.13 $\pm$ 0.75 77.23 $\pm$ 0.24 80.21 $\pm$ 0.11 81.25 $\pm$ 0.10 81.23 $\pm$ 0.14
Table 14: Experiment setting III: Main results on CMU-MOSEI
Test Modality Dataset: CMU-MOSEI / Evaluation Metric: F1
Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
40.75 $\pm$ 0.14 56.02 $\pm$ 0.56 31.38 $\pm$ 0.61 58.97 $\pm$ 0.53 55.85 $\pm$ 0.69 53.59 $\pm$ 1.07 60.10 $\pm$ 0.76 61.56 $\pm$ 0.22 61.78 $\pm$ 0.58
38.15 $\pm$ 0.92 37.12 $\pm$ 0.66 31.73 $\pm$ 0.16 38.26 $\pm$ 2.40 29.40 $\pm$ 1.67 33.29 $\pm$ 2.89 36.28 $\pm$ 3.51 39.76 $\pm$ 0.23 39.36 $\pm$ 1.45
40.82 $\pm$ 0.12 56.06 $\pm$ 0.29 31.86 $\pm$ 0.23 58.58 $\pm$ 0.31 57.02 $\pm$ 1.60 53.38 $\pm$ 0.65 58.70 $\pm$ 2.10 61.37 $\pm$ 0.28 61.15 $\pm$ 0.39
40.37 $\pm$ 0.46 55.70 $\pm$ 0.35 31.59 $\pm$ 0.23 61.20 $\pm$ 0.46 54.02 $\pm$ 2.88 53.08 $\pm$ 0.40 58.58 $\pm$ 0.71 61.84 $\pm$ 0.70 61.78 $\pm$ 0.58
42.16 $\pm$ 0.22 55.89 $\pm$ 0.37 30.84 $\pm$ 0.26 59.35 $\pm$ 0.12 55.87 $\pm$ 0.68 54.11 $\pm$ 0.49 58.42 $\pm$ 0.27 61.56 $\pm$ 0.61 61.94 $\pm$ 0.37
42.45 $\pm$ 0.85 56.06 $\pm$ 0.29 30.71 $\pm$ 0.71 60.00 $\pm$ 0.85 56.64 $\pm$ 1.65 52.92 $\pm$ 1.11 59.50 $\pm$ 0.80 61.10 $\pm$ 0.10 61.00 $\pm$ 0.76
Appendix F Visualization of Experiment
In this section, we provide visual evidence to further support our claims regarding expert selection dynamics and modality robustness.
Figure 4 shows the UMAP projection of multi-modal embeddings on CMU-MOSI. It clearly demonstrates that imputed modalities form distributions consistent with their original counterparts, suggesting that our imputation module maintains semantic alignment across modalities.
Figure 4: UMAP Visualization of Multi-modal Embedding Space: CMU-MOSI
Figure 5 illustrates the softmax gating mechanism without regularization, which exhibits severe expert collapse where only a few experts are consistently selected throughout training. In contrast, Figure 6 shows that our proposed ConfNet gating maintains both load balance and expert specialization, effectively mitigating collapse and enabling more stable and diverse expert routing.
Figure 5: Softmax Score-based Expert Selection: Serious Expert Collapse when training goes on
Figure 6: ConfNet Gating Expert Selection: Expert collapse is consistently mitigated throughout training. While certain experts are preferred, less frequently selected experts are still actively utilized. Moreover, expert selection becomes increasingly balanced as training progresses, indicating improved stability and diversity in routing.
Figures 7 to 10 compare various gating mechanisms under similar conditions.
Figure 7: Softmax Gate with Load Balance Loss Expert Selection: An optimization conflict arises, often reflected in a "Sharp Sinusoidal Wave" pattern, experts that were heavily selected in one epoch tend to be suppressed in the next. This oscillatory behavior indicates that the typical load balancing loss does not promote true specialization or expertise among experts. Instead, it merely enforces uniform usage, potentially at the cost of performance, by prioritizing balanced selection over actual expert quality.
Figure 8: Laplacian Gating Expert Selection: Laplacian gating can partially alleviate the load imbalance issue and achieves better expert utilization than softmax. However, as training progresses, selection becomes increasingly skewed, with only a few experts being consistently chosen. This eventually leads to expert collapse, where most experts become inactive and underutilized.
Figure 9: Mean Gating Expert Selection: The "Sharp Sinusoidal Wave" pattern persists under mean gating, where experts heavily selected in one epoch are suppressed in the next. This oscillatory behavior reflects ambiguity in expert selection, stemming from the absence of explicit guidance in the routing mechanism. While mean gating achieves better load balance than Softmax with $\mathcal{L}_{\text{load}}$, it ultimately results in inferior performance, as shown in Table 4.
Figure 10: Gaussian Gating Expert Selection: The "Sharp Sinusoidal Wave" pattern persists under Gaussian gating, where experts heavily selected in one epoch are suppressed in the next. This oscillatory behavior reflects ambiguity in expert selection, stemming from the inappropriate gating score guided from Gaussian. While mean gating achieves better load balance than Softmax with $\mathcal{L}_{\text{load}}$, it ultimately results in inferior performance, as shown in Table 4.
Finally, Figures 11 and 12 report Experiment III results. ConfMoE demonstrates superior resilience across increasing missing rates, with both F1 and AUC metrics degrading gracefully. These results visually confirm the effectiveness of confidence-guided gating and modality imputation under challenging conditions.
Figure 11: Experiment III: CMU-MOSEI Performance Plot
Figure 12: Experiment III: CMU-MOSI Performance Plot
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: [N/A]
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: [N/A]
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: [N/A]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: [N/A]
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: [N/A]
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: [N/A]
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: [N/A]
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: [N/A]
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: [N/A]
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: [N/A]
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: [N/A]
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: [N/A]
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: [N/A]
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: [N/A]
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.

### v2

@@
 However, existing SMoE approach often lacks proper ability to handle missing modality, leading to performance degradation and poor generalization in real-world applications.
-We propose Conf-SMoE to introduce a two-stage imputation module to handle the missing modality problem for the SMoE architecture and reveal the insight of expert collapse from theoretical analysis with strong empirical evidence.
-Inspired by our theoretical analysis, Conf-SMoE propose a novel expert gating mechanism by detaching the softmax routing score to task confidence score w.r.t ground truth.
+We propose ConfSMoE to introduce a two-stage imputation module to handle the missing modality problem for the SMoE architecture by taking the opinion of experts and reveal the insight of expert collapse from theoretical analysis with strong empirical evidence.
+Inspired by our theoretical analysis, ConfSMoE propose a novel expert gating mechanism by detaching the softmax routing score to task confidence score w.r.t ground truth signal.
 This naturally relieves expert collapse without introducing additional load balance loss function.
 We show that the insights of expert collapse aligns with other gating mechanism such as Gaussian and Laplacian gate.
-We also evaluate the proposed method on four different real world dataset with three different experiment settings to conduct comprehensive the analysis of Conf-SMoE on modality fusion and resistance to missing modality.
+The proposed method is evaluated on four different real world dataset with three distinct experiment settings to conduct comprehensive analysis of ConfSMoE on resistance to missing modality and the impacts of proposed gating mechanism.
 We release our official implementation in: https://github.com/IcurasLW/Official-Repository-of-ConfSMoE.git
 1 Introduction
-As modern applications increasingly involve data from multiple heterogeneous sources, multimodal learning has emerged as a critical paradigm to unlock complementary insights and enhance decision-making in complex tasks spanning medical decision support[39, 41], cross-modal inference[31, 5], temporal reasoning[40, 23], and natural language understanding[9, 21].
-Sparse Mixture-of-Experts (SMoE) architectures[34, 35, 22] are widely used in multimodal learning for their ability to assign different experts to different input patterns through conditional computation.
+As modern applications increasingly involve data from multiple heterogeneous sources, multimodal learning has emerged as a critical paradigm to unlock complementary insights and enhance decision-making in complex tasks spanning medical decision supportZheng et al.
+(2024a, c), cross-modal inferenceWu et al.
+(2017); Cheng et al.
+(2024), temporal reasoningZheng et al.
+(2024b); Ni et al.
+(2024), and natural language understandingFedus et al.
+(2022); Masoudnia and Ebrahimpour (2014).
+Sparse Mixture-of-Experts (SMoE) architecturesYao et al.
+(2024); Yun et al.
+(2024); Mustafa et al.
+(2022) are widely used in multimodal learning for their ability to assign different experts to different input patterns through conditional computation.
 They are particularly effective when all input modalities are present, enabling diverse feature extraction and flexible expert routing.
@@
 To reduce this complexity, recent SMoE frameworks typically activate the Top-K experts per sample.
-However, the use of softmax-based routing[9] tends to produce sharp distributions, often concentrating the routing on a single expert.
-This leads to the expert collapse problem [9, 12, 35], where only a few experts dominate the computation as shown in Figure 1(a), resulting in poor diversity and limited generalization.
+However, the use of softmax-based routingFedus et al.
+(2022) tends to produce sharp distributions, often concentrating the routing on a single expert.
+This leads to the expert collapse problem Fedus et al.
+(2022); Han et al.
+(2024); Yun et al.
+(2024), where only a few experts dominate the computation as shown in Figure 1(a), resulting in poor diversity and limited generalization.
 Efforts to mitigate this issue using load-balancing losses introduce additional optimization challenges.
@@
 Figure 1: Expert Collapse of MoE and Comparison with Two Different Modality Imputation
-While recent works[9, 12, 35] has tackled expert collapse within SMoE architectures, handling missing modalities remains an orthogonal and unresolved challenge.
-In the context of missing modalities, many recent approaches aim to reconstruct absent views through cross-modal information sharing [20, 29, 34] as shown in Figure 1(b).
+While recent worksFedus et al.
+(2022); Han et al.
+(2024); Yun et al.
+(2024) has tackled expert collapse within SMoE architectures, handling missing modalities remains an orthogonal and unresolved challenge.
+In the context of missing modalities, many recent approaches aim to reconstruct absent views through cross-modal information sharing Ma et al.
+(2021); Wang et al.
+(2023); Yao et al.
+(2024) as shown in Figure 1(b).
 However, these methods typically introduce multiple auxiliary losses to jointly learn task-relevant, shared, and modality-specific features, which increases training complexity and may inadvertently erode modality-unique information.
-To alleviate this issue, recent work has proposed using learnable modality banks[35] to impute missing inputs based on the available modality combinations.
-Nevertheless, such banks are usually initialized randomly, similar to prompt tuning[15, 18], and often fail to capture fine-grained modality-specific structure, limiting their effectiveness in practical settings.
-To overcome the limited robustness of existing SMoE models under incomplete modality conditions, we propose Confidence-guided Sparse Mixture-Of-E (Conf-SMoE), a unified framework that introduces two novel mechanisms for handling arbitrary missing modalities.
+To alleviate this issue, recent work has proposed using learnable modality banksYun et al.
+(2024) to impute missing inputs based on the available modality combinations.
+Nevertheless, such banks are usually initialized randomly, similar to prompt tuningJia et al.
+(2022); Lester et al.
+(2021), and often fail to capture fine-grained modality-specific structure, limiting their effectiveness in practical settings.
+To overcome the limited robustness of existing SMoE models under incomplete modality conditions, we propose Confidence-guided Sparse Mixture-Of-E (ConfSMoE), a unified framework that introduces two novel mechanisms for handling arbitrary missing modalities.
 Our method stems from a core insight: modality-specific structure should be preserved and leveraged for both reconstruction and routing, a property often overlooked in prior work.
 Based on this, we propose a two-stage imputation strategy that first estimates missing modalities using intra-modality correlations, and then refines these estimates through instance-level cross-modal alignment, as illustrated in Figure 1(c).
-The rationale behind this approach is inspired by the modality gap phenomenon discussed by [19], which suggests that a certain level of discrepancy between modalities is both expected and beneficial for downstream tasks.
+The rationale behind this approach is inspired by the modality gap phenomenon discussed by Liang et al.
+(2022), which suggests that a certain level of discrepancy between modalities is both expected and beneficial for downstream tasks.
 Thus, it is natural to first impute missing modalities using intra-modality information subsequently enhance them through instance-level cross-modal refinement.
-We further uncover and formalize a critical cause of expert collapse in SMoE: the gradient concentration induced by softmax gating, which suppresses routing diversity and limits generalization[9, 30, 35, 12, 6].
+We further uncover and formalize a critical cause of expert collapse in SMoE: the gradient concentration induced by softmax gating, which suppresses routing diversity and limits generalizationFedus et al.
+(2022); Wang et al.
+(2024); Yun et al.
+(2024); Han et al.
+(2024); Chi et al.
+(2022).
 Through theoretical analysis, we show that conventional load-balancing losses exacerbate this issue by introducing gradient conflicts during optimization, reflecting by the "Sinoual Wave" pattern in expert selection plot shown in Figure 3(a).
@@
 This formulation not only avoids the adverse effects of entropy-based balancing but also enables stable, interpretable expert selection.
-By decoupling routing from the optimization bottlenecks of prior designs, Conf-SMoE achieves both enhanced expert diversity and consistent performance across a wide range of incomplete modality scenarios.
+By decoupling routing from the optimization bottlenecks of prior designs, ConfSMoE achieves both enhanced expert diversity and consistent performance across a wide range of incomplete modality scenarios.
 Our contributions are the following:
@@
 In this work, we build upon the SMoE framework by incorporating expert routing into the Transformer’s feedforward layers, enabling dynamic token-to-expert assignment based on learned routing scores.
-This architecture supports selective expert activation, where only the Top-K experts are engaged per token, and proves especially effective in the joint expert setting[12], allowing tokens from different modalities to be routed to shared experts for enhanced cross-modal interaction.
-We begin with the standard Softmax SMoEs as a baseline configuration.
-Denote that input token embedding $\mathbf{h}\in\mathbb{R}^{d}$ and the a learnable softmax router $G(\mathbf{h})=softmax(\mathbf{h})$ that produces routing score $g_{i}$ for each expert $E_{i}$.
-Then, Top-K function is applied to select the top-k expert in forward, the final tokens/samples representation is weighted summation from Top-K expert with corresponding routing score $g_{i}$.
+This architecture supports selective expert activation, where only the Top-K experts are engaged per token, and proves especially effective in the joint expert setting Han et al.
+(2024), allowing tokens from different modalities to be routed to shared experts for enhanced cross-modal interaction.
+We begin with the standard Softmax SMoEs with $N$ number of expert as a baseline configuration.
+Denote that input token embedding $\mathbf{h}\in\mathbb{R}^{d}$ and the a learnable softmax router $G(\mathbf{h})=softmax(\mathbf{h})$ that produces routing score $g_{i}$ for each expert $E_{i}$ from expert pool $\mathbf{E}=[E_{1},E_{2},\dots,E_{N}]\in\mathbb{R}^{d\times d}$.
+Then, Top-K function is applied to select the Top-K experts in forward, the final tokens/samples representation is weighted summation from Top-K experts with corresponding routing score $g_{i}$.
 Formally, we can represent the forward process of SMoE as:
-$E(\mathbf{h})=\mathbf{h}+\sum_{i=1}^{K}G(\mathbf{h})E_{i}(\mathbf{h})$ (1)
-Where $E_{i}(\cdot)$ represent the i-th expert chosen by the router in expert pool and we would like to sum up the representation from Top-K expert by gating score out of $G(\mathbf{h})$.
+$E(\mathbf{h})=\mathbf{h}+\sum_{i}^{K}g_{i}E_{i}(\mathbf{h})$ (1)
 2.2 What Causes Expert Collapse?
-The Jacobian matrix of Eq. 1 with respect to $\mathbf{h}$ can be analyzed to reveal the underlying cause of expert collapse in SMoE. Therefore,
-$\mathbf{J}_{MoE}=\sum_{i=1}^{K}\frac{\partial softmax(\mathbf{h})}{\partial\mathbf{h}}E_{i}(\mathbf{h})+\sum_{i=1}^{K}G(\mathbf{h})E_{i}^{{}^{\prime}}(\mathbf{h})$ (2)
-Let’s denote $\mathbf{g}=softmax(\mathbf{u})$, where $\mathbf{u}\in\mathbb{R}^{N}$ is the gating logits to select expert.
-The Jacobian matrix for the softmax function is denoted as $\mathbf{J}_{softmax}\in\mathbb{R}^{d\times d}$, where each element $\mathbf{J}_{ij}=\frac{\partial softmax(\mathbf{h_{j})}}{\partial\mathbf{h_{j}}}=g_{i}(\delta_{ij}-g_{i})$ and $\delta_{ij}$ is Kronecker delta.
-In matrix form, we can express Jacobian of softmax as: $\mathbf{J}_{softmax}=\textbf{diag(g)}-\textbf{g}\textbf{g}^{T}$, and the set of expert forward function $\mathbf{E}_{K}(\mathbf{h})=[E_{1}(\mathbf{h}),E_{2}(\mathbf{h}),\dots,E_{K}(\mathbf{h})]\in\mathbb{R}^{d\times d}$.
-For simplicity, we simplify the Top-K operation into $\mathbf{E_{K}}$.
-Therefore, we can rewrite Eq.
-2 to Eq.
-3: the first term is to learn a better routing score driven by gradient of gating score $\mathbf{g}$ and second term is to learn a better representation as the gradient driven by $E_{i}^{{}^{\prime}}(\mathbf{h})$.
-The Top-k mechanism makes the first term of Eq.
-3 sparse and only a few experts contribute to gradient while only one expert dominate the gradient due to softmax sharp distribution.
-It is clear that the greatest gating score from the sharp distribution of softmax dominates the gradient in training phrase, resulting expert collapse based on token preference and produce rich-get-richer expert.
-$\mathbf{J}_{MoE}=\underbrace{\mathbf{E}_{K}(\mathbf{h})(\textbf{ diag(g) }-\textbf{g}\textbf{g}^{\top})}_{\text{Learning a better routing score}}\;\;\;\;\;+\underbrace{\sum_{i=1}^{K}G(\mathbf{h})E_{i}^{{}^{\prime}}(\mathbf{h})}_{\text{Learning a better representation}}$ (3)
+The Jacobian matrix of Eq. 1 w.r.t $\mathbf{h}$ can be analyzed to reveal the underlying cause of expert collapse in SMoE. Therefore,
+$\mathbf{J}_{MoE}=\sum_{i}^{K}\frac{\partial g_{i}}{\partial\mathbf{h}}E_{i}(\mathbf{h})+\sum_{i}^{K}g_{i}E_{i}^{{}^{\prime}}(\mathbf{h})$ (2)
+[…]
+$\mathbf{J}_{MoE}=\underbrace{\mathbf{E}_{K}(\mathbf{h})(\textbf{diag(g)}_{K}-\textbf{g}\textbf{g}_{K}^{\top})}_{\text{Learning a better routing score}}\;\;\;\;\;+\underbrace{\sum_{i}^{K}g_{i}E_{i}^{{}^{\prime}}(\mathbf{h})}_{\text{Learning a better representation}}$ (3)
 2.3 Gradient of Load Balance Loss
 This expert collapse phenomenon restrict the model learning capacity since many other expert are wasted and cannot reflect the actual expert diversity.
-Two widely-used solutions to achieve balance expert usage is to introduce additional load balance loss and expert capacity.
-However, such load balance loss approach is difficult to optimize since it produces gradient direction opposite to softmax routing score terms in Eq 3.
-We consider that any load balance loss $\mathcal{L}_{load}$ will eventually results a flat and diverse softmax distribution, which encourage higher entropy on the softmax distribution.
+One widely-used solutions to achieve balance expert usage is to introduce additional load balance loss Yun et al.
+(2024); Han et al.
+(2024).
+However, such load balance loss approach is difficult to optimize since it produces gradient direction opposite to routing score terms in Eq 3.
+To show this, we consider that any load balance loss $\mathcal{L}_{load}$ will eventually results a diverse softmax routing distribution, which encourage higher entropy on the routing score distribution.
 From Assumption 1 defined in Appendix, we define that the objective of any $\mathcal{L}_{load}$ as $\frac{1}{\mathcal{H(\mathbf{g})}}$ over the softmax distribution, where $\mathcal{H(\mathbf{g})}=-\sum_{i=1}^{N}g_{i}log\ g_{i}$ represent the entropy of softmax distribution g.
@@
 Recall that $\mathbf{E}_{K}(\mathbf{h})$ in Eq.
-3 is a set of expert with ReLU activation, we assume this network should not output negative value which is theoretically true in most cases.
-We know the fact that the Jacobian of softmax $\mathbf{J}_{softmax}$ is symmetry and always positive semi-definite[10], making the optimization of better routing score is always positive.
-In contrast, $log(\textbf{g})+1=[log(g_{1})+1,log(g_{2})+1,\dots,log(g_{K})+1]$.
-When sharp distribution appears, only one $g_{j}$ is closed to 1 while other $g_{i},(i\neq j)$ are closed to 0, which is practically and mostly true.
-Therefore, $\frac{1}{\mathcal{H(\textbf{g})}^{2}}(log(\textbf{g})+1)^{\top}$ is exponentially dominated by negative value and $\mathbf{J}_{load}$ is opposite to the first term of $\mathbf{J}_{MoE}$.
+3 is a set of expert with ReLU activation, we assume this network should not output negative value as activation is located on positive side.
+We also know the fact that the Jacobian of softmax $\mathbf{J}_{softmax}$ is symmetry and always Positive Semi-Definite (PSD) Gao and Pavel (2017), making the optimization of better routing score is always positive.
+In contrast, $log(\textbf{g})+1=[log(g_{1})+1,log(g_{2})+1,\dots,log(g_{N})+1]$.
+When sharp distribution appears, only one $g_{j}$ is closed to 1 while other $g_{i},(i\neq j)$, are closed to 0.
+Therefore, $\frac{1}{\mathcal{H(\textbf{g})}^{2}}(log(\textbf{g})+1)^{\top}$ is exponentially dominated by negative value, producing conflict gradient between $\mathbf{J}_{load}$ and the first term of $\mathbf{J}_{MoE}$.
+The Top-K operation does not change PSD property of term, but the gradient will be backward through the off Top-K term in $J_{MoE}$.
+It is also worthy to notice that gradient in Eq.
+4 conflict occurs for all Top-K number as the load balance loss consider all $g_{i}$ regardless of number K expert selection.
+Taking an extreme example, Top-1 selection Fedus et al.
+(2022) would generate one $g_{i}$ out of g for the second term of Eq.
+3 while Eq.
+4 still take all $g_{i}$ to gradient computation.
 Therefore, the opposite gradient from axillary load balance loss makes the training difficult to converge to global minimum and Expert selection becomes ambiguous, failing to reflect actual capacity of expert.
@@
 4 and Eq.
-3, we understand that load balance loss shows opposite gradient direction to softmax router.
-Then, the potential solution to explicitly achieve load balance is to remove the association of sharp distribution from Eq.
-3.
+3, we understand that load balance loss shows opposite gradient direction to softmax router in typical SMoE.
+To avoid this gradient conflict and balance the load, the potential solution is to remove the association of sharp distribution from Eq.
+3 and attempt to explicitly achieve load balance.
 One is to replace the routing function as illustrated in Figure.
@@
 2 (d).
-A typical example from a recent work, FuseMoE [12], they replace the Softmax router to Laplacian router and calculate the routing score based on similarity of the expert embedding $W_{N}$ by $\mathcal{L}_{1}$ distance.
-Then the Top-K experts are selected with highest similarity to expert embedding.
+A typical example from a recent work, FuseMoE Han et al.
+(2024), they replace the Softmax router to Laplacian router and calculate the routing score based on similarity of the expert embedding by $\mathcal{L}_{1}$ distance.
+Then the Top-K experts are selected with highest similarity to expert embedding and Laplacian router is not a sharp distribution so that the load is explicit balanced.
 The second approach is to decouple the connection of gating score and expert selection.
@@
 Therefore, we propose a Confidence-Guided Gating Network Pool, which introduces a novel routing strategy guided by task confidence.
-Specifically, instead of solely relying on conventional softmax-based gating, which often results in sharp distributions and entangled gradient contributions, our method leverages the ground-truth label information to supervise the routing mechanism, encouraging expert selection aligned with task-relevant confidence rather than token or embedding similarity.
-For each expert $E_{i}$, we introduce an auxiliary confidence estimation network: $U_{i}$ that maps token embedding $\mathbf{h}\in\mathbb{R}^{d}$ to a single scalar logits $v_{i}\in\mathbb{R}$.
-$v_{i}$ is then mapped to a probability $g_{i}\in(0,1)$ as gating score by Sigmoid, representing the confidence of ground truth label $y_{t}$.
+Specifically, instead of solely relying on conventional softmax-based gating, which often results in sharp distributions and entangled gradient contributions, our method leverages the ground-truth label information to supervise the routing mechanism, encouraging expert selection aligned with task-relevant confidence rather than token or embedding similarity as in Han et al.
+(2024); Fedus et al.
+(2022).
+For each expert $E_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$, we introduce an auxiliary confidence estimation network: a one-layer linear network $U_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$ that maps token embedding $\mathbf{h}\in\mathbb{R}^{d}$ to a single scalar logits $v_{i}\in\mathbb{R}^{1}$.
+$v_{i}$ is then mapped to a confidence score $c_{i}\in(0,1)$ as gating score by Sigmoid, representing the confidence of ground truth label $y_{t}$.
 To learn the true confidence $p_{t}$ of downstream task w.r.t $y_{t}$, where $t$ is the notation for ground truth.
-We apply MSE loss as an axillary loss that minimize the difference between $g_{i}$ and $p_{t}$.
+We apply MSE loss as an axillary loss that minimize the difference between $c_{i}$ and $p_{t}$.
 Formally, given any classification dataset $D$ with supervised signal $y_{i}$ and number of instance $|D|$, we consider the task loss $\mathcal{L}_{task}$ as Cross Entropy Loss and the final objective is to optimize:
-$\mathcal{L}=\mathcal{L}_{task}+\mathcal{L}_{conf}=\frac{1}{|D|}\sum^{D}p_{t}log(p_{t})+\frac{1}{|D|K}\sum^{D}\sum^{K}_{i=1}(g_{i}-p_{t})^{2}$ (5)
-The intuition is that the learning of gating score in Eq.
-3 will be driven by the task-direct score, this forces model to select expert with task-preference instead of token preference as Softmax gating [9] or expert embedding preference as Laplacian gating does [12].
-During inference, we use the learned confidence score $g_{i}$ from the ConfNet as a confidence proxy for the task-specific relevance of each expert.
-These scores are used for specialized token fusion, enabling dynamic integration of modality-specific and task-relevant information.
-The $g_{i}$ is Token-level confidence, which each token will have one corresponding $g_{i}$ for aggregation.
-Additionally, to capture coarse-grained signals, we aggregate token-level representations assigned to each expert and average them to compute expert-level gating signals as Expert-level variant, denoted as ConfSMoE-E in experiment.
+$\mathcal{L}=\mathcal{L}_{task}+\mathcal{L}_{conf}=\frac{1}{|D|}\sum^{D}y_{t}log(p_{t})+\frac{1}{|D|K}\sum^{D}\sum^{K}_{i=1}(c_{i}-p_{t})^{2}$ (5)
+Note that Eq.
+5 is convex as both cross-entropy and MSE loss are well-known convex.
+During both training and inference, $c_{i}$ serves as a proxy gating score for each expert to token fusion in Eq.
+2.
+The $c_{i}$ is Token-level confidence, which each token has one corresponding $c_{i}$ for aggregation as in $\sum^{K}_{i}G(\mathbf{h})E_{i}(\mathbf{h})$ as in Eq.
+1, resulting in a variant of proposed method for fine-grained feature learning, denoted as ConfSMoE-T.
+Additionally, a expert-level gating signal is also implemented for coarse-grained information retrieval.
+We aggregate token-level representations assigned to each expert and average them to compute expert-level gating signals as Expert-level variant, denoted as ConfSMoE-E in experiment.
+The multimodal interaction is learned during the token assignment for all experts, which all modalities are shared one identical expert pool.
 This confidence-driven routing enables both fine-grained specialization and robust multimodal interaction, leading to improved generalization and interpretability.
-Figure 2: The Framework of Conf-SMoE: (a) Conf-SMoE consider the confidence learned from ConfNet pool as the gating score for expert aggregation.
+Figure 2: The Framework of ConfSMoE: (a) ConfSMoE consider the confidence learned from ConfNet pool as the gating score for expert aggregation.
+The ConfNet pool is a set of one-layer linear weight $\mathbf{U}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{1}$.
 Missing modality imputation is achieved by taking the mean of random instances from the corresponding modality of training data as common modality specific feature.
 The instance-specific feature are sparsely aggregated from existing modality.
-(b) Expert collapse is because of the softmax sharp gating score.
+(b) Sharp gating score causes expert collapse.
 (c)(d) "Fair" gating score distribution and decoupling gating score can alleviate the expert collapse without the helps of load balance loss.
 3.2 Missing Modality Modeling: Two-Stage Imputation
-Inspired by the modality gap phenomenon observed in contrastive learning [19], we hypothesize that a similar gap exists in our multimodal setting—and that it may even contribute positively to model performance, as illustrated in Appendix Figure 4.
-The presence of this modality gap suggests that imputing a missing modality using only shared information from the remaining modalities is inherently challenging, due to the distinct representational spaces each modality occupies.
+Inspired by the modality gap phenomenon observed in contrastive learning Liang et al.
+(2022), we hypothesize that a similar gap exists in our multimodal setting, and that it may even contribute positively to model performance, as illustrated in Appendix Figure 4.
+The presence of this modality gap suggests that imputing a missing modality using only shared information from the available modalities is inherently challenging, due to the distinct representational spaces each modality occupies.
 Ideally, the imputed modality should preserve the characteristics of its original distribution without being significantly biased by other modalities.
-To address this, we propose a Two-stage Imputation strategy as following:
-Pre-Imputation: When a modality $M_{i}\in\mathbb{R}^{d}$ is missing, we generate an pre-imputed modality embedding $\bar{M}_{i}$ by sampling $K$ instances from the modality pool $M_{i,n}=[M_{i,o},M_{i,1},\cdots,M_{i,n}]$, where $M_{i,n}$ represent the $i$-th modality from $n$-th instance modality pool.
-The modality pool contains all existing modality data for given number of modalities in training set.
-To remove the instance-specific feature and preserve the modality-common feature, we take an average over the $K$ instances as pre-imputed feature.
-The randomized selection of instances from the training set further aids in eliminating instance-specific noise during backpropagation.
-Thus, we send the existing and pre-imputed modality into backbone model to extract interaction between modality.
-Naturally, increasing $K$ yields more representative modality features but reduces stochasticity, potentially causing all imputed modalities to converge to the same representation across different missing-modality samples and the interaction between other modalities may be more deterministic and uninformative.
-We consider this $K$ as one of hyperparameter to finetune.
-$\bar{M}_{i}=\frac{1}{K}\sum^{K}_{n=1}\mathbf{1}M_{i,n}$ (6)
-Where $\mathbf{1}\in\{0,1\}^{n}$ is the selection vector for $M_{i}$ and $\mathbf{1}_{i}=1$ represents the $i$-th instance from modality pool is selected.
-Post-Imputation: To add up instance-specific information from other modality, we propose a post-modality imputation to further capture the interactions after the MoE forward.
-Unlike conventional approaches that solely rely on the fused representation from one network, our method leverages information from multiple experts networks by selectively incorporating tokens beyond the Top-K experts.
-This allows the model to absorb nuanced, instance-specific features and interaction between modalities from different views of expert.
-To implement this, we employ a sparse attention mechanism, under the assumption that only a subset of modality-specific information is meaningfully shared across existing modalities.
-By enforcing sparsity, the model is encouraged to attend only to relevant parts of the available modalities, thereby reducing the risk of injecting modality-irrelevant or noisy information into the imputed representation, especially in heterogeneous or weakly correlated multimodal settings.
-Specifically, given a missing modality $M_{i}\in\mathbb{R}^{s\times d}$ and existing modality $M_{j}\in\mathbb{R}^{s\times d}$, where $s$ is the sequence length and $d$ is the hidden dimension.
-Each token of existing modality $t_{s}\in\mathbb{R}^{d}$ is sent to Top-K different experts to generate $K$ specialized tokens $t_{s,e}\in\mathbb{R}^{d}$, where $s$ represents the sequential position and $e$ represents expert index.
-We concatenate the specialized tokens $t_{s,e}$ in sequence dimension and apply cross attention operation over those specialized tokens and select only Top-K correlated specialized tokens for $\bar{M}_{i}$ to generate final imputed modality $M_{i}^{*}$.
-$M_{i}^{*}=\bar{M}_{i}+SparseCrossAttention(\bar{M}_{i},M_{j},M_{j})$ (7)
-We set the number of tokens $K$ selected from $t_{s,e}$ as $K=\frac{s(|M|-1)}{B}$, where $|M|$ represents the number of modality in dataset and $B$ is a sparsity hyperparameter, we set $B$ as 4 by default.
+To address this, we propose a Two-stage Imputation strategy to reconstruct the representative feature from the missing modality distribution itself, and further refine the imputed modality by adding up instance-specific feature from the available modalities.
+Pre-Imputation: To retrieve modality-specific feature, given a sample $x_{i}$ with missing modality $M_{m}\in\mathbb{R}^{s\times d}$ and arbitrary available modality $M_{a}\in\mathbb{R}^{s\times d}$, where $s$ is the sequence length.
+Pre-imputation aims to obtain the modality-specific feature from other instances of $M_{m}$ distribution.
+One typical way is to take the mean feature distribution of $M_{m}$, but simply taking the mean may be overfitting to training set and all representation across different modality samples will be deterministic and uninformative.
+In contrast to mean representation, we wish to introduce some stochasticity during the pre-imputation as this stochasticity may prevent the representation to collapse to the same representation.
+Specifically, we generate an pre-imputed modality embedding $\bar{M}_{m}$ by sampling $n$ number of instances out of the modality pool $M_{m,i}=[M_{m,1},M_{m,2},\cdots,M_{m,n}]$.
+$\bar{M}_{m}=\frac{1}{n}\sum^{n}_{i=1}\mathbf{1}M_{m,i}$ (6)
+Where $\mathbf{1}\in\{0,1\}^{n}$ is the selection vector for $M_{m}$ and $\mathbf{1}_{i}=1$ represents the $i$-th instance from modality pool is selected.
+$\bar{M}_{m}$ will be sent to ConfSMoE for learning the interaction across different modalities.
+We denoted $M_{m}^{*}$ as the pre-imputed modality representation after ConfSMoE and $M^{k}_{a}$ as the available modality representation after k-th expert.
+Post-Imputation: To refine $M^{*}_{m}$ by available modality, we proposed a post-modality imputation to further capture the interactions between available and missing modality.
+Unlike conventional approaches that solely rely on the shared representation from one network, our method leverages information from Top-K experts networks by selectively incorporating tokens.
+This allows the model to absorb nuanced, instance-specific features and interaction between modalities from different views of expert to available modality.
+We employ a sparse attention mechanism, under the assumption that only a subset of modality-specific information is meaningfully shared across $M^{k}_{a}$.
+By enforcing sparsity, the model is encouraged to attend to the most relevant parts of the available modalities, thereby reducing the risk of injecting modality-irrelevant or noisy information into the imputed representation.
+Recall from Eq.
+1, Top-K MoE generate $K$ specialized representation from available modality $M_{a}^{k}\in\mathbb{R}^{s\times d}$, $k=0,1,\dots K$.
+Then, we concatenate all $M_{a}^{k}$ returned from $K$ expert to form opinion representation from different expert, denoted as $M_{a}^{*}\in\mathbb{R}^{(s\times K)\times d}$.
+We apply a sparse cross attention over $M_{m}^{*}$ and $M_{a}^{*}$ to select most relevant tokens to refine $M_{m}^{*}\in\mathbb{R}^{s\times d}$.
+The SparseCrossAttention takes $M_{m}^{*}$ as Query modality $\mathbf{Q}=\mathbf{W}_{q}M_{m}^{*}$, $M_{a}^{*}$ as Key modality $\mathbf{K}=\mathbf{W}_{k}M_{a}^{*}$ and Value modality $\mathbf{V}=\mathbf{W}_{v}M_{a}^{*}$, where $\mathbf{Q}\in\mathbb{R}^{s\times d}$ and $\mathbf{K},\mathbf{V}\in\mathbb{R}^{(s\times K)\times d}$.
+Note that $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are shared between all modalities.
+Then, we obtained the cross attention score $\mathbf{A}\in\mathbb{R}^{s\times(s\times K)}$ from the dot product of $\mathbf{Q}$ and $\mathbf{K}$.
+Then, we apply a binary mask $\mathbf{M}$ over the attention scores $\mathbf{A}$ to retain only the Top-T entry based on the rank of attention score (1 indicate presence) alongside the dimension of $\mathbf{K}$, yielding a sparse attention map $\mathbf{A}^{*}$.
+$T=\frac{s(|M|-1)}{B}$, where $|M|$ represents the number of modality in dataset and $B$ is a sparsity hyperparameter, we set $B$ as 4 by default.
+$\mathbf{A}^{*}=\mathbf{M}\odot softmax(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}),\in\mathbb{R}^{s\times(s\times K)}$ (7)
+Thus, we aggregate the value matrix $V$ using the sparse attention weights $A^{*}$ to obtain the output of $SparseCrossAttention$ (SCA).
+When multiple available modality $M_{a}^{*}$ are presented in the dataset, we compute their contributions individually and sum them to refine $M_{m}^{*}$:
+$M^{*}_{m}=LayerNorm\left(M_{m}^{*}+\sum^{a}SCA(M_{m}^{*},M_{a}^{*},M_{a}^{*})\right)=LayerNorm\left(M_{m}^{*}+\sum^{a}\mathbf{A}^{*}\mathbf{V}\right)$ (8)
+In addition, we expect this refinement to redistribute the representation of $M_{m}^{*}$ instead of generating large scale representation and lead to diverge training.
+A LayerNorm layer is applied to scale the feature map back to consistent distribution as other layers.
+With SCA, the Post-imputation serves as a refinement module to add up instance-specific feature on $M_{m}^{*}$, which each modality can contribute differently on different sub-segment representation of $M_{m}^{*}$.
+We visualize the attention weights after imputation in the Appendix.
+Figures 13, 14, and 15 illustrate these weights, showing that the query modality can dynamically adjust the significance of information by leveraging both expert specialization and shared cross-modal cues.
+The attention maps reveal distinct weight distributions across different modalities and experts.
 4 Experiment
-4.1 Dataset
-MIMIC-III [16] is a public dataset contains irregular time series and clinical notes modality.
-We follow the preprocessing scripts from MMIMI-III benchmark [13] and [38] to form multimodal patient instance.
-However, [38] filter the instances with missing modality and we modified the scripts to preserve with instances with missing modality.
-MIMIC-IV [17] is a public dataset contains irregular time series, clinical notes, ECG signal and chest image.
-We follow the preprocessing scripts from MedFuse [14] to preprocess irregular time series and further preprocess clinical notes, ECG signal and chest image by ourself.
-We follow the dataset split in MedFuse [14] and form multimodal patient instance.
-CMU-MOSI [36] is a multimodal sentiment analysis dataset comprising audio, video, and text modalities.
-CMU-MOSEI [2] is an extension of the CMU-MOSI dataset, utilizing the same modalities—audio, text, and video from YouTube recordings.
-We follow preprocessing script of CMU-MOSI and CMU-MOSEI from MMML[22].
-4.2 Baselines
-SMIL[20] impute missing modality with Bayesian meta-learning.
-ShaSpec[29] leverages multiple loss function to learn the modality-share and modality-specific information and impute the missing modality with modality-shared information.
-TF[37] extract the modality interaction by multimodal sub-networks and tensor fusion layer.
-mmFormer[28] is a typical multimodal transformer with multihead attention mechanism.
-LlMoE[22] achieves the multimodal learning with contrastive learning.
-FuseMoE[12] respect the missing modality and jointly learn single and missing modality combination data input and FuseMoE implement Laplacian gating mechanism for modality fusoin and expert selection.
-In experiment, we denote FuseMoE-S as softmax gating and FuseMoE-L as Laplacian gating.
-FlexMoE[35] assigns the combination of modality to different experts and impute the missing modality by modality bank.
-For our approach, we use ConfSMoE-T denotes Token-Level confidence fusion and ConfSMoE-E denotes Expert-level confidence fusion.
 4.1 Experiment Setting
@@
 To eliminate confounding effects introduced by varying encoder architectures, we use the same modality-specific encoders for all models, as differences in encoder design can significantly influence final performance.
-For the MIMIC-III and MIMIC-IV datasets, we employ a pretrained ClinicalBERT[1] for textual data, a one-layer patch embedding module for both irregular time series and ECG signals, and a pretrained DenseNet[7] from TorchXRayVision for chest X-ray (CXR) images.
+For the MIMIC-III and MIMIC-IV datasets, we employ a pretrained ClinicalBERT Alsentzer et al.
+(2019) for textual data, a one-layer patch embedding module for both irregular time series and ECG signals, and a pretrained DenseNet Cohen et al.
+(2022) from TorchXRayVision for chest X-ray (CXR) images.
 Experiment Setting I: Natural Missing Modality.
@@
 Table 1: Experiment setting I: Main results on MIMIC-IV
-Task Metric SMIL ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
+Task Metric SMIL ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
 48-IHM F1 39.58 $\pm$ 1.12 32.97 $\pm$ 1.08 45.39 $\pm$ 1.60 11.60 $\pm$ 0.54 43.76 $\pm$ 0.44 30.14 $\pm$ 1.15 40.21 $\pm$ 1.29 35.29 $\pm$ 0.30 49.18 $\pm$ 0.22 48.32 $\pm$ 0.30
@@
 AUC 63.90 $\pm$ 1.31 62.57 $\pm$ 0.23 70.33 $\pm$ 0.12 56.28 $\pm$ 0.16 72.50 $\pm$ 0.67 55.48 $\pm$ 0.39 58.52 $\pm$ 0.27 71.57 $\pm$ 0.47 74.56 $\pm$ 0.19 74.42 $\pm$ 0.34
-Table 2: Experiment setting II: Main results on CMU-MOSI
-Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
-0% F1 41.78 $\pm$ 0.64 43.28 $\pm$ 0.36 50.22 $\pm$ 0.20 50.91 $\pm$ 0.19 47.69 $\pm$ 0.39 48.15 $\pm$ 0.57 51.50 $\pm$ 0.55 51.83 $\pm$ 0.27 51.70 $\pm$ 0.29
-AUC 65.08 $\pm$ 0.39 64.48 $\pm$ 0.51 74.62 $\pm$ 0.49 78.12 $\pm$ 0.31 75.26 $\pm$ 0.20 75.96 $\pm$ 0.29 79.38 $\pm$ 0.23 80.56 $\pm$ 0.10 80.62 $\pm$ 0.07
-10% F1 43.40 $\pm$ 0.74 40.14 $\pm$ 0.13 48.66 $\pm$ 1.20 49.91 $\pm$ 0.33 46.81 $\pm$ 0.45 48.81 $\pm$ 0.42 49.92 $\pm$ 0.28 49.99 $\pm$ 0.85 50.89 $\pm$ 0.07
-AUC 65.45 $\pm$ 0.56 61.54 $\pm$ 0.66 77.21 $\pm$ 0.76 77.12 $\pm$ 0.14 70.25 $\pm$ 0.93 72.36 $\pm$ 0.50 77.07 $\pm$ 0.21 78.06 $\pm$ 0.64 78.03 $\pm$ 0.21
-20% F1 39.15 $\pm$ 0.13 42.64 $\pm$ 0.30 48.51 $\pm$ 1.15 48.84 $\pm$ 0.61 47.92 $\pm$ 0.58 45.00 $\pm$ 0.44 49.16 $\pm$ 0.24 50.58 $\pm$ 0.73 50.17 $\pm$ 0.31
-AUC 60.87 $\pm$ 0.29 65.52 $\pm$ 0.98 74.16 $\pm$ 0.71 76.64 $\pm$ 0.23 71.05 $\pm$ 0.36 72.15 $\pm$ 0.37 73.80 $\pm$ 1.43 77.17 $\pm$ 0.55 77.18 $\pm$ 0.11
-30% F1 39.54 $\pm$ 0.17 40.42 $\pm$ 0.75 46.66 $\pm$ 1.24 46.13 $\pm$ 0.55 43.78 $\pm$ 0.51 44.48 $\pm$ 0.63 45.49 $\pm$ 0.44 47.57 $\pm$ 0.10 47.52 $\pm$ 0.77
-AUC 60.80 $\pm$ 0.44 60.57 $\pm$ 0.99 71.25 $\pm$ 0.57 74.33 $\pm$ 0.49 70.38 $\pm$ 0.56 72.25 $\pm$ 0.48 72.54 $\pm$ 0.92 75.59 $\pm$ 0.18 75.42 $\pm$ 0.49
-40% F1 35.32 $\pm$ 0.32 38.69 $\pm$ 0.42 42.22 $\pm$ 0.44 43.31 $\pm$ 0.38 41.88 $\pm$ 0.42 43.20 $\pm$ 0.20 43.13 $\pm$ 0.61 45.12 $\pm$ 0.27 44.18 $\pm$ 0.10
-AUC 59.58 $\pm$ 0.13 61.62 $\pm$ 0.07 68.44 $\pm$ 0.14 67.21 $\pm$ 0.43 64.61 $\pm$ 0.63 66.14 $\pm$ 0.11 68.12 $\pm$ 0.34 71.65 $\pm$ 0.47 69.55 $\pm$ 0.49
-50% F1 36.39 $\pm$ 0.97 38.46 $\pm$ 1.09 41.14 $\pm$ 0.31 41.54 $\pm$ 0.77 38.55 $\pm$ 0.34 40.90 $\pm$ 0.49 41.95 $\pm$ 0.48 43.89 $\pm$ 0.51 42.48 $\pm$ 0.12
-AUC 55.60 $\pm$ 0.12 60.49 $\pm$ 0.80 61.87 $\pm$ 0.67 67.43 $\pm$ 0.34 62.28 $\pm$ 0.28 65.73 $\pm$ 0.30 66.47 $\pm$ 0.49 71.51 $\pm$ 0.53 69.04 $\pm$ 0.32
-Experiment Setting II: In Table 2 and Appendix E Table 10, we implement Random Modality Dropout to study the sensitivity of model to missing modality.
+Experiment Setting II: In Appendix Table 9 and Table 10, we implement Random Modality Dropout to study the sensitivity of model to missing modality.
 In this setting, we empirically find that those methods with missing modality imputation usually shows higher tenacity in high proportion of missingness such as FlexMoE and our propose method.
@@
 Our proposed method, on the other hand, shows superior performance across all setting in CMU-MOSEI and CMU-MOSI for all metrics.
-Experiment Setting III: In Appendix E Table 11, Table 12, Table 13 and Table 14, we study which modality combination is significant and how is the tenacity of our proposed method to different distribution shift on missing modality.
+Experiment Setting III: In Appendix Table 11, Table 12, Table 13 and Table 14, we study which modality combination is significant and how is the tenacity of our proposed method to different distribution shift on missing modality.
 We find that single modality shows very similar performance during inference phrase in our proposed method and Flex-MoE across AUC and F1, indicating the effectiveness of missing modality imputation.
@@
 We show some visualization evidence in Figure 4 that the embedding of those two modalities are mostly symmetry.
-Table 3: Ablation I: Module dropout. CMU-MOSI adopt 50% missing in Experiment III
-MIMIC-III CMU-MOSI
-F1 AUC F1 AUC
-w/o impute 43.79 $\pm$ 2.81 85.33 $\pm$ 1.22 42.51 $\pm$ 2.25 69.67 $\pm$ 1.27
-w/o Conf 48.59 $\pm$ 1.97 85.91 $\pm$ 0.84 42.47 $\pm$ 0.94 68.77 $\pm$ 1.66
-w/o impute & Conf 46.32 $\pm$ 1.66 85.15 $\pm$ 0.57 42.50 $\pm$ 0.63 66.73 $\pm$ 1.41
-Conf-SMoE 53.44 $\pm$ 0.27 87.05 $\pm$ 0.58 43.89 $\pm$ 0.51 71.51 $\pm$ 0.53
-4.5 Are Missing Modality Imputation and ConfNet Actually Useful?
-We further implement Ablation Study I in Table 3 by dropping off the module we designed in proposed method and study which part of of model is actually useful.
-Specifically, we dropout the modality imputation module and keep the missing modality as zeros, we can see that F1 and AUC dropped off by 6.21% and 1.81% on MIMIC-III, 5.62% and 0.81% on CMU-MOSI.
-This performance drops is even worse when imputation and ConfNet gating mechanism is removed from proposed method, providing strong empirical evidence for effectiveness of our design.
-It is interesting that the w/o impute & Conf variant outperform w/o impute, indicating the model may only focus on existing modality instead of modality interaction.
 4.3 Why Does Confidence-Guided Gating Improve Performance?
@@
 We also found that considering the uniform weight for all experts can achieve similar performance as sole softmax and generate similar sinusoidal pattern as softmax w/ $\mathcal{L}_{load}$, shown in Figure 9.
-(a) Expert selection of softmax w/ $\mathcal{L}_{laod}$
+(a) Expert selection of softmax w/ $\mathcal{L}_{load}$
 (b) Expert selection of ConfNet
 Figure 3: Expert selection: X-axis represents epoch and Y-axis represents expert ID. The darker the color, the more selection will be for a particular expert. See detailed and more plots in Appendix F
-Table 4: Ablation II: Different router. CMU-MOSI adopt 50% missing setting in Experiment III
+Table 2: Ablation I: Different router. CMU-MOSI adopt 50% missing setting in Experiment III
 Task Metric Mean Softmax Softmax w/ $\mathcal{L}_{load}$ Gaussian Laplacian ConfNet
@@
 AUC 85.13 $\pm$ 0.16 85.91 $\pm$ 0.84 85.97 $\pm$ 0.52 86.70 $\pm$ 0.33 85.47 $\pm$ 1.05 87.05 $\pm$ 0.58
-CMU-MOSI F1 42.29 $\pm$ 0.87 41.47 $\pm$ 0.94 42.43 $\pm$ 1.34 42.57 $\pm$ 1.81 43.15 $\pm$ 1.53 43.89 $\pm$ 0.51
-AUC 68.01 $\pm$ 1.68 68.77 $\pm$ 1.66 67.40 $\pm$ 0.54 67.74 $\pm$ 0.45 67.60 $\pm$ 1.54 71.51 $\pm$ 0.53
+CMU-MOSI F1 42.29 $\pm$ 0.87 41.47 $\pm$ 0.94 42.43 $\pm$ 1.34 42.57 $\pm$ 1.81 43.15 $\pm$ 1.53 44.34 $\pm$ 0.77
+AUC 68.01 $\pm$ 1.68 68.77 $\pm$ 1.66 67.40 $\pm$ 0.54 67.74 $\pm$ 0.45 67.60 $\pm$ 1.54 70.41 $\pm$ 1.31
 In addition, Laplacian and Gaussian gating function can achieve comparable to softmax w/ $\mathcal{L}_{load}$ gating mechanism without helps of $\mathcal{L}_{load}$, providing further evidence to support the advantage of balance expert selection.
@@
 Notably, both ConfNet and Laplacian gates demonstrate a common pattern: they maintain the specialization of dominant experts while redistributing some load to less active ones, which highlights that strict load balancing may suppress expert specialization and degrade performance, whereas maintaining a balance between specialization and load distribution leads to superior results.
+4.4 How ConfNet and Modality Imputation integrate together?
+Table 3: Ablation II: Module dropout. CMU-MOSI adopt 50% missing in Experiment II
+Variants MIMIC-III CMU-MOSI
+F1 AUC F1 AUC
+w/o Conf 48.59 $\pm$ 1.97 85.91 $\pm$ 0.84 42.47 $\pm$ 0.94 68.77 $\pm$ 1.66
+w/o impute 49.59 $\pm$ 0.29 86.31 $\pm$ 0.12 42.46 $\pm$ 2.29 68.44 $\pm$ 1.82
+w/o post-impute 49.86 $\pm$ 1.89 86.53 $\pm$ 0.10 43.14 $\pm$ 1.32 69.13 $\pm$ 1.45
+w/o impute & Conf 46.32 $\pm$ 1.66 85.15 $\pm$ 0.57 41.98 $\pm$ 0.74 68.48 $\pm$ 1.90
+ConfSMoE-T 53.44 $\pm$ 0.27 87.05 $\pm$ 0.58 44.34 $\pm$ 0.77 70.41 $\pm$ 1.13
+We implement the experiment in Table 3 by removing ConfNet and the two-stage imputation module gradually.
+This experiment demonstrates that both novel modules, ConfNet and the two-stage imputation, are effective on their own, while their integration provides strong robustness against the missing modality challenge.
+Specifically, ConfNet captures informative features from the imputed modality, and the adaptive two-stage imputation operation reconstructs high-quality features, thereby complementing each other to achieve robust performance.
+Specifically, we dropped the entire imputation module and keep the missing modality as zeros (w/o impute), we can see that F1 and AUC dropped off by 3.85 and 0.74 on MIMIC-III, 1.88 and 1.97 on CMU-MOSI.
+This performance drops is even worse when imputation and ConfNet gating mechanism is removed from proposed method, providing strong empirical evidence for effectiveness of our design.
+We further remove the post-imputation (w/o post-impute) and observe that ConfNet can still take advantage of modality-specific feature for their downstream task, showing robustness against missing modality.
 4.5 Complexity Analysis
-Table 5: Complexity analysis on CMU-MOSI with 50% missing setting in Experiment III
-Metrics ShaSpec TF LlMoE mmFormer FuseMoE FlexMoE ConfSMoE
-MFLOPs 0.51 8.42 1016.15 1577.63 199.80 20.16 47.10
-#Params 102,022 4,211,331 3,457,819 11,373,315 243,543,180 1,126,147 3,116,310
-F1 score 36.39 $\pm$ 0.97 38.46 $\pm$ 1.09 41.54 $\pm$ 0.77 41.14 $\pm$ 0.31 40.90 $\pm$ 0.49 41.95 $\pm$ 0.30 43.89 $\pm$ 0.51
-In Table 5, we compared the complexity between our proposed method and baseline models in terms of FLOPs and number of parameters.
-ShaSpec is the simplest models, but it turns out not comparable performance in many tasks.
-On the other hand, those MoE-based approach like FuseMoE and LlMoE are extremely complicated either in FLOPs and number of parameters while the performance is suboptimal compare to FlexMoE and our proposed method.
-mmFormer show very comparable results with our proposed method, however, the FLOPs and parameters are extremely larger than ConfSMoE, leading to poor training and inference efficiency.
-On the other hand, FlexMoE is 2 times less complicated than ConfSMoE while the performance is worse than LlMoE, mmFormer and ConfSMoE.
-By balancing complexity and performance from Table 5, it is clear that ConfSMoE is the best model to balance complexity and performance.
+In Table 4, we control the number of parameters for both baseline and proposed methods to compare model complexity and performance.
+Although ShaSpec and TF achieve lower FLOPs for the same parameter amount, their performance remains less competitive than most Transformer- and MoE-based models, except FuseMoE.
+Among MoE-based models, ConfSMoE-T delivers the best performance while maintaining similar MFLOPs and parameter counts to the second-best baseline.
+Under comparable parameter settings, LIMoE and mmFormer fail to keep FLOPs as low as ConfSMoE-T and yield suboptimal results.
+Moreover, LIMoE is particularly inefficient, as its MFLOPs are substantially higher than those of other models with similar parameter sizes.
+Overall, ConfSMoE-T achieves the best balance between efficiency and performance among MoE-based baselines.
+Table 4: Complexity analysis I: Small Scale
+Metrics ShaSpec TF LlMoE mmFormer FuseMoE FlexMoE ConfSMoE-T
+MFLOPs 21,73 4.20 1245.03 237.39 143.93 40.32 39,62
+#Params 2,177,406 2,101,571 2,130,227 2,378,371 20,971,020 2,184,963 2,184,730
+F1 score 39.12 $\pm$ 1.21 35.81 $\pm$ 0.85 41.11 $\pm$ 0.90 41.14 $\pm$ 0.31 36.52 $\pm$ 0.84 42.12 $\pm$ 1.18 44.63 $\pm$ 1.65
 5 Related Works
 While in many works considers that the interaction between multimodal data can bring comprehensive understanding for downstream task.
-Initial multimodal fusion approach usually incoperates neural kernel fusion[4, 24], early/late network fusion [32, 3, 11].
-Approaches such as Tensor Fusion Network[37], Multimodal Transformer[28] and Multimodal Adaption Gate[25], LlMoE[22] all highlights the multimodal interactions are beneficial to downstream task.
-In medical domain, MedFuse [14] incorporated CXR and EHR modality to jointly fuse modality with early and later fusion in a single fusion network.
-MISTS [38] incorporated multi-time attention mechanism to encoder irregular time serires and adapts multimodal cross attention to fuse two different modality.
-[33] fused multimodal data in a attention gate and compute a replacement vector for modality fusion.
+Initial multimodal fusion approach usually incoperates neural kernel fusionBucak et al.
+(2013); Poria et al.
+(2015), early/late network fusion Xu et al.
+(2023); Baltrušaitis et al.
+(2018); Guo et al.
+(2019).
+Approaches such as Tensor Fusion NetworkZadeh et al.
+(2017), Multimodal TransformerTsai et al.
+(2019) and Multimodal Adaption GateRahman et al.
+(2020), LlMoEMustafa et al.
+(2022) all highlights the multimodal interactions are beneficial to downstream task.
+In medical domain, MedFuse Hayat et al.
+(2022) incorporated CXR and EHR modality to jointly fuse modality with early and later fusion in a single fusion network.
+MISTS Zhang et al.
+(2023) incorporated multi-time attention mechanism to encoder irregular time serires and adapts multimodal cross attention to fuse two different modality.
+Yang and Wu (2021) fused multimodal data in a attention gate and compute a replacement vector for modality fusion.
 However, these approaches are limited to full modality only and difficult to extend to increasing number of modality since the complexity can be significantly increased by the number of modality.
 Missing modality is realistic and practical problem in real life application such as medical diagnose.
-SMIL [20] proposed a Bayesian meta-learning solution with reconstruction network to impute the missing modality from in hidden state during fusion.
-[8, 26] propose generative adversarial networks to impute the missing modality by semi-supervised learning.
-ShaSpec [29] and DrFuse [34] designed multiple loss function to learn share and specific representation of different modality.
+SMIL Ma et al.
+(2021) proposed a Bayesian meta-learning solution with reconstruction network to impute the missing modality from in hidden state during fusion.
+Du et al.
+(2018); Shang et al.
+(2017) propose generative adversarial networks to impute the missing modality by semi-supervised learning.
+ShaSpec Wang et al.
+(2023) and DrFuse Yao et al.
+(2024) designed multiple loss function to learn share and specific representation of different modality.
 However, the reconstruction network from SMIL and ShaSpec requires multiple additional auxiliary loss function, which occupies the great amount of gradient in optimization process instead of task specific loss.
-[27] In addition, recent studies like FuseMoE [12] respect the missingness of modalities and wish to assign each combination of multimodal data to different experts.
-Flex-MoE[35] further expand the advantages of FuseMoE and introduce a modality bank to impute the missing modality.
+Sun et al.
+(2024) In addition, recent studies like FuseMoE Han et al.
+(2024) respect the missingness of modalities and wish to assign each combination of multimodal data to different experts.
+Flex-MoEYun et al.
+(2024) further expand the advantages of FuseMoE and introduce a modality bank to impute the missing modality.
 6 Conclusion
-In this work, a novel Conf-SMoE model, a confidence-guided sparse mixture-of-experts framework, is proposed to robustly handle multimodal learning with missing modalities.
+In this work, a novel ConfSMoE model, a confidence-guided sparse mixture-of-experts framework, is proposed to robustly handle multimodal learning with missing modalities.
 By theoretically analyzing the root cause of expert collapse and empirically validating our findings, we propose two key innovations: a two-stage imputation strategy that preserves modality-specific structure, and a novel confidence-driven gating mechanism that decouples expert selection from softmax-induced sharpness.
@@
 Appendix A Notation
-To facilitate understanding of the model formulation and derivations, Table 6 outlines the notations used throughout this paper.
-These include variables associated with expert routing, gating mechanisms, modality assignments, and sparsity control, which are central to our proposed Conf-SMoE framework.
 Table 5: Notation
 Variable Definition
-$E_{i}$ i-th Expert Forward Function
+$E_{i}$ i-th Expert Function
+E Expert pool
+$\textbf{E}_{K}$ Set of Top-K expert
+$\mathbf{U}$ ConfNet pool
 $\mathbf{h}$ Embedding of Arbitrary Data Input
 $K$ Number of K selection
-$N$ Number of Expert
+$T$ Number of T selection in Post-Imputation
+NN Total number of Expert
+$n$ Number of instances selected in Pre-imputation
 $G(\cdot)$ Gating Function
@@
 $\delta_{ij}$ Kronecker delta
-$u_{i}$ softmax input at entry i
+$u_{i}$ Softmax input at i-th entry
+$c_{i}$ Confidence score for Expert $E_{i}$
+$g_{i}$ Gating score for Expert $E_{i}$
+To facilitate understanding of the model formulation and derivations, Table 5 outlines the notations used throughout this paper.
+These include variables associated with expert routing, gating mechanisms, modality assignments, and sparsity control, which are central to our proposed ConfSMoE framework.
 Appendix B Derivation of Jacobian for MoE and Load Balance Loss
 B.1 Jacobian for Softmax
-This section provides detailed mathematical derivations of the Jacobian matrix used in the gating function of Conf-SMoE and its relation to the load balance loss.
+This section provides detailed mathematical derivations of the Jacobian matrix used in the gating function of ConfSMoE and its relation to the load balance loss.
 We first derive the Jacobian of the softmax operation applied to the gating logits, which is essential for analyzing expert selection dynamics.
@@
 where $u_{i}$ is the i-th input value of gate logits.
-$\frac{\partial\textbf{g}}{\partial u}=\left\{\begin{aligned} \frac{e^{u_{i}}(\sum^{N}_{j=1}e^{u_{j}}-e^{u_{i}})}{(\sum^{N}_{j=1}e^{u_{j}})^{2}}&\quad\text{if }i=j\\ -\frac{e^{u_{i}}e^{u_{j}}}{(\sum^{N}_{j=1}e^{u_{j}})^{2}}&\quad\text{if }i\neq j\end{aligned}\right.$ (9)
+$\frac{\partial\textbf{g}}{\partial u}=\left\{\begin{aligned} \frac{e^{u_{i}}(\sum^{N}_{j=1}e^{u_{j}}-e^{u_{i}})}{(\sum^{N}_{j=1}e^{u_{j}})^{2}}&\quad\text{if }i=j\\ -\frac{e^{u_{i}}e^{u_{j}}}{(\sum^{N}_{j=1}e^{u_{j}})^{2}}&\quad\text{if }i\neq j\end{aligned}\right$.
+(10)
 $\frac{\partial\textbf{g}}{\partial u}=\left\{\begin{aligned} g_{i}(1-g_{j})&\quad\text{if }i=j\\ -g_{i}g_{j}&\quad\text{if }i\neq j\end{aligned}\right$. (11)
@@
 Appendix C Experiment Details
-C.1 Dataset and Task Description
+C.1 Dataset
+MIMIC-III Johnson et al.
+[2016] is a public dataset contains irregular time series and clinical notes modality.
+We follow the preprocessing scripts from MMIMI-III benchmark Harutyunyan et al.
+[2019] and Zhang et al.
+[2023] to form multimodal patient instance.
+However, Zhang et al.
+[2023] filter the instances with missing modality and we modified the scripts to preserve with instances with missing modality.
+MIMIC-IV Johnson et al.
+[2023] is a public dataset contains irregular time series, clinical notes, ECG signal and chest image.
+We follow the preprocessing scripts from MedFuse Hayat et al.
+[2022] to preprocess irregular time series and further preprocess clinical notes, ECG signal and chest image by ourself.
+We follow the dataset split in MedFuse Hayat et al.
+[2022] and form multimodal patient instance.
+CMU-MOSI Zadeh et al.
+[2016] is a multimodal sentiment analysis dataset comprising audio, video, and text modalities.
+CMU-MOSEI Bagher Zadeh et al.
+[2018] is an extension of the CMU-MOSI dataset, utilizing the same modalities—audio, text, and video from YouTube recordings.
+We follow preprocessing script of CMU-MOSI and CMU-MOSEI from MMMLMustafa et al.
+[2022].
+C.2 Baselines
+SMILMa et al.
+[2021] impute missing modality with Bayesian meta-learning.
+ShaSpecWang et al.
+[2023] leverages multiple loss function to learn the modality-share and modality-specific information and impute the missing modality with modality-shared information.
+TFZadeh et al.
+[2017] extract the modality interaction by multimodal sub-networks and tensor fusion layer.
+mmFormerTsai et al.
+[2019] is a typical multimodal transformer with multihead attention mechanism.
+LlMoEMustafa et al.
+[2022] achieves the multimodal learning with contrastive learning.
+FuseMoEHan et al.
+[2024] respect the missing modality and jointly learn single and missing modality combination data input and FuseMoE implement Laplacian gating mechanism for modality fusoin and expert selection.
+In experiment, we denote FuseMoE-S as softmax gating and FuseMoE-L as Laplacian gating.
+FlexMoEYun et al.
+[2024] assigns the combination of modality to different experts and impute the missing modality by modality bank.
+For our approach, we use ConfSMoE-T denotes Token-Level confidence fusion and ConfSMoE-E denotes Expert-level confidence fusion.
+C.3 Dataset and Task Description
 Table 6: Dataset statistics
@@
 LOS 6621 Time Series, Text 2 Yes 24.21%
-25-PHE 12,278 Time Series, Text 2 Yes 24.25%
-MIMIC-IV 48-IHM 22,733 Time Series, Text 2 Yes 47.10%
-LOS 22,733 Time Series, Text 2 Yes 47.10%
-25-PHE 47,027 Time Series, Text 2 Yes 50.32%
+25-PHE 12,278 Time Series, Text 25 Yes 24.25%
+MIMIC-IV 48-IHM 22,733 Image, ECG, Time Series, Text 2 Yes 47.10%
+LOS 22,733 Image, ECG, Time Series, Text 2 Yes 47.10%
+25-PHE 47,027 Image, ECG, Time Series, Text 25 Yes 50.32%
 CMU-MOSI Sentiment Analysis 1870 Audio, Video, Text 3 Yes 0% - 50%
@@
 We attempt to predict one 25 acute care conditions presented in a givem ICU stay record.
-The definition of those 25 acute care conditions are defined in MIMIC benchmark [16]
+The definition of those 25 acute care conditions are defined in MIMIC benchmark Johnson et al.
+[2016]
 CMU-MOSI and CMU-MOSEI are multimodal sentiment analysis benchmarks consisting of video clips annotated with sentiment scores.
@@
   • MIMIC-IV download link: https://physionet.org/content/mimiciv/2.2/ and benchmark: https://github.com/nyuad-cai/MedFuse
-C.2 Hyperparameter and Hardware Setting
-All experiments are conducted in a remote server with 4 $\times$ RTX4090 24GB with 256 RAM and 96-cores CPU.
+C.4 Hyperparameter and Hardware Setting
+All experiments are conducted in a remote server with 4 $\times$ RTX4090 24 GB with 256 GB RAM and 96-cores CPU.
 Table 7: Hyperparameter setting
@@
 #Expert 8 for MIMIC-III and MIMIC-IV, 4 for CMU-MOSI and CMU-MOSEI
-Top-K Expert 2
+#Top-K Expert 2
 Learning Rate 3e-4
@@
 Weight for $\mathcal{L}_{conf}$ 1
-# K for Pre-imputation 10
+#Top-T for Pre-imputation 10
 Random Seed 2023, 2024, 2025
@@
 Appendix D Limitation
-This work may be limited to the task with supervision signal and additional RAM memory usage is required since the modality pool requires additional space for saving modality data.
+This work may be limited to expert discontinuity as it is build upon on sparse MoE backbone.
+However, the trade-off between sparse MoE and continuous / fully differentiable MoE, more computational cost vs continuity of experts, is still underexplored.
 Appendix E Additional Results
@@
 LOS F1 62.01 $\pm$ 0.54 57.27 $\pm$ 0.36 61.10 $\pm$ 0.65 56.23 $\pm$ 0.79 62.75 $\pm$ 0.10 62.26 $\pm$ 0.36 62.27 $\pm$ 1.61 64.18 $\pm$ 0.34 66.27 $\pm$ 0.40 66.39 $\pm$ 0.23
-AUC 80.08 $\pm$ 0.35 74.71 $\pm$ 0.07 77.94 $\pm$ 0.12 72.60 $\pm$ 0.79 79.80 $\pm$ 0.27 80.05 $\pm$ 0.51 79.70 $\pm$ 0.43 80.49 $\pm$ 0.48 81.66$\pm$ 0.26 81.71 $\pm$ 0.18
+[…]
 25-PHE F1 33.01 $\pm$ 0.30 19.52 $\pm$ 0.39 35.10 $\pm$ 0.38 30.38 $\pm$ 0.38 36.22 $\pm$ 0.39 33.48 $\pm$ 0.22 34.07 $\pm$ 0.51 35.31 $\pm$ 0.57 37.40 $\pm$ 0.13 40.13 $\pm$ 0.40
 AUC 75.17 $\pm$ 1.84 64.61 $\pm$ 0.25 71.79 $\pm$ 0.26 66.56 $\pm$ 0.17 77.93 $\pm$ 0.44 77.22 $\pm$ 0.29 77.19 $\pm$ 0.33 77.37 $\pm$ 0.10 78.93 $\pm$ 0.38 78.28 $\pm$ 0.73
+Table 9: Experiment setting II: Main results on CMU-MOSI
+Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
+0% F1 41.78 $\pm$ 0.64 43.28 $\pm$ 0.36 50.22 $\pm$ 0.20 50.91 $\pm$ 0.19 47.69 $\pm$ 0.39 48.15 $\pm$ 0.57 51.50 $\pm$ 0.55 51.83 $\pm$ 0.27 51.70 $\pm$ 0.29
+AUC 65.08 $\pm$ 0.39 64.48 $\pm$ 0.51 74.62 $\pm$ 0.49 78.12 $\pm$ 0.31 75.26 $\pm$ 0.20 75.96 $\pm$ 0.29 79.38 $\pm$ 0.23 80.56 $\pm$ 0.10 80.62 $\pm$ 0.07
+10% F1 43.40 $\pm$ 0.74 40.14 $\pm$ 0.13 48.66 $\pm$ 1.20 49.91 $\pm$ 0.33 46.81 $\pm$ 0.45 48.81 $\pm$ 0.42 49.92 $\pm$ 0.28 49.99 $\pm$ 0.85 50.89 $\pm$ 0.07
+AUC 65.45 $\pm$ 0.56 61.54 $\pm$ 0.66 77.21 $\pm$ 0.76 77.12 $\pm$ 0.14 70.25 $\pm$ 0.93 72.36 $\pm$ 0.50 77.07 $\pm$ 0.21 78.06 $\pm$ 0.64 78.03 $\pm$ 0.21
+20% F1 39.15 $\pm$ 0.13 42.64 $\pm$ 0.30 48.51 $\pm$ 1.15 48.84 $\pm$ 0.61 47.92 $\pm$ 0.58 45.00 $\pm$ 0.44 49.16 $\pm$ 0.24 50.58 $\pm$ 0.73 50.17 $\pm$ 0.31
+AUC 60.87 $\pm$ 0.29 65.52 $\pm$ 0.98 74.16 $\pm$ 0.71 76.64 $\pm$ 0.23 71.05 $\pm$ 0.36 72.15 $\pm$ 0.37 73.80 $\pm$ 1.43 77.17 $\pm$ 0.55 77.18 $\pm$ 0.11
+30% F1 39.54 $\pm$ 0.17 40.42 $\pm$ 0.75 46.66 $\pm$ 1.24 46.13 $\pm$ 0.55 43.78 $\pm$ 0.51 44.48 $\pm$ 0.63 45.49 $\pm$ 0.44 47.57 $\pm$ 0.10 47.52 $\pm$ 0.77
+AUC 60.80 $\pm$ 0.44 60.57 $\pm$ 0.99 71.25 $\pm$ 0.57 74.33 $\pm$ 0.49 70.38 $\pm$ 0.56 72.25 $\pm$ 0.48 72.54 $\pm$ 0.92 75.59 $\pm$ 0.18 75.42 $\pm$ 0.49
+40% F1 35.32 $\pm$ 0.32 38.69 $\pm$ 0.42 42.22 $\pm$ 0.44 43.31 $\pm$ 0.38 41.88 $\pm$ 0.42 43.20 $\pm$ 0.20 43.13 $\pm$ 0.61 45.12 $\pm$ 0.27 44.18 $\pm$ 0.10
+AUC 59.58 $\pm$ 0.13 61.62 $\pm$ 0.07 68.44 $\pm$ 0.14 67.21 $\pm$ 0.43 64.61 $\pm$ 0.63 66.14 $\pm$ 0.11 68.12 $\pm$ 0.34 71.65 $\pm$ 0.47 69.55 $\pm$ 0.49
+50% F1 36.39 $\pm$ 0.97 38.46 $\pm$ 1.09 41.14 $\pm$ 0.31 41.54 $\pm$ 0.77 38.55 $\pm$ 0.34 40.90 $\pm$ 0.49 41.95 $\pm$ 0.48 44.34 $\pm$ 0.77 42.48 $\pm$ 0.12
+AUC 55.60 $\pm$ 0.12 60.49 $\pm$ 0.80 61.87 $\pm$ 0.67 67.43 $\pm$ 0.34 62.28 $\pm$ 0.28 65.73 $\pm$ 0.30 66.47 $\pm$ 0.49 70.41 $\pm$ 1.31 69.04 $\pm$ 0.32
 Table 10: Experiment setting II: Main results on CMU-MOSEI
-Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
+Missing Rate Metric ShaSpec TF mmFormer LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
 0% F1 47.48 $\pm$ 0.27 38.89 $\pm$ 0.54 56.75 $\pm$ 0.47 59.13 $\pm$ 0.52 57.04 $\pm$ 0.15 58.71 $\pm$ 0.23 60.62 $\pm$ 0.96 61.31 $\pm$ 0.14 62.35 $\pm$ 0.12
@@
 AUC 65.32 $\pm$ 0.59 60.09 $\pm$ 0.35 69.31 $\pm$ 0.10 73.24 $\pm$ 0.76 69.85 $\pm$ 0.11 70.05 $\pm$ 0.46 72.08 $\pm$ 1.27 73.63 $\pm$ 0.48 73.88 $\pm$ 0.25
-50% F1 42.75 $\pm$ 1.22 33.01 $\pm$ 0.51 47.39 $\pm$ 0.16 46.69 $\pm$ 0.63 43.08 $\pm$ 0.47 $4$$\pm$ 0.36 46.19 $\pm$ 0.34 47.31 $\pm$ 0.20 48.13 $\pm$ 0.33
+50% F1 42.75 $\pm$ 1.22 33.01 $\pm$ 0.51 47.39 $\pm$ 0.16 46.69 $\pm$ 0.63 43.08 $\pm$ 0.47 44.44 $\pm$ 0.36 46.19 $\pm$ 0.34 47.31 $\pm$ 0.20 48.13 $\pm$ 0.33
 AUC 62.32 $\pm$ 0.42 58.20 $\pm$ 0.17 66.61 $\pm$ 0.22 70.90 $\pm$ 0.82 67.63 $\pm$ 0.55 67.44 $\pm$ 0.16 70.15 $\pm$ 1.10 71.74 $\pm$ 0.14 70.48 $\pm$ 0.11
@@
 Test Modality Dataset: CMU-MOSI / Evaluation Metric: AUC
-Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
+Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
 62.18 $\pm$ 0.88 72.72 $\pm$ 0.08 59.37 $\pm$ 1.12 74.63 $\pm$ 0.13 55.35 $\pm$ 1.11 56.44 $\pm$ 1.48 74.59 $\pm$ 1.50 75.57 $\pm$ 0.55 75.56 $\pm$ 0.33
@@
 Test Modality Dataset: CMU-MOSI / Evaluation Metric: F1
-Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
+Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
 41.27 $\pm$ 0.81 48.25 $\pm$ 0.11 34.96 $\pm$ 0.72 44.45 $\pm$ 0.33 30.51 $\pm$ 2.53 29.90 $\pm$ 2.58 50.54 $\pm$ 0.73 54.07 $\pm$ 0.29 54.04 $\pm$ 0.36
@@
 Test Modality Dataset: CMU-MOSEI / Evaluation Metric: AUC
-Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
+Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
 64.47 $\pm$ 0.10 76.60 $\pm$ 0.39 56.98 $\pm$ 0.20 81.23 $\pm$ 0.07 75.81 $\pm$ 0.82 76.62 $\pm$ 0.37 80.63 $\pm$ 0.41 81.45 $\pm$ 0.02 81.52 $\pm$ 0.07
@@
 Test Modality Dataset: CMU-MOSEI / Evaluation Metric: F1
-Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfMoE-T ConfMoE-E
+Video Text Audio ShaSpec mmFormer TF LlMoE FuseMoE-S FuseMoE-L FlexMoE ConfSMoE-T ConfSMoE-E
 40.75 $\pm$ 0.14 56.02 $\pm$ 0.56 31.38 $\pm$ 0.61 58.97 $\pm$ 0.53 55.85 $\pm$ 0.69 53.59 $\pm$ 1.07 60.10 $\pm$ 0.76 61.56 $\pm$ 0.22 61.78 $\pm$ 0.58
@@
 Figure 12: Experiment III: CMU-MOSI Performance Plot
-NeurIPS Paper Checklist
-  1. 1.
-    Claims
-  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
-  3. Answer: [Yes]
-  4. Justification: [N/A]
-  5. Guidelines:
-    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
-    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
-A No or NA answer to this question will not be perceived well by the reviewers.
-    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
-    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
-  6. 2.
-    Limitations
-  7. Question: Does the paper discuss the limitations of the work performed by the authors?
-  8. Answer: [Yes]
-  9. Justification: [N/A]
-  10. Guidelines:
-    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
-    • The authors are encouraged to create a separate "Limitations" section in their paper.
-    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
-The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
-    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
-In general, empirical results often depend on implicit assumptions, which should be articulated.
-    • The authors should reflect on the factors that influence the performance of the approach.
-For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
-Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
-    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
-    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
-    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
-The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
-Reviewers will be specifically instructed to not penalize honesty concerning limitations.
-  11. 3.
-    Theory assumptions and proofs
-  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
-  13. Answer: [Yes]
-  14. Justification: [N/A]
-  15. Guidelines:
-    • The answer NA means that the paper does not include theoretical results.
-    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
-    • All assumptions should be clearly stated or referenced in the statement of any theorems.
-    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
-    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
-    • Theorems and Lemmas that the proof relies upon should be properly referenced.
-  16. 4.
-    Experimental result reproducibility
-  17.
-Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
-  18. Answer: [Yes]
-  19. Justification: [N/A]
-  20. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
-    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
-    • Depending on the contribution, reproducibility can be accomplished in various ways.
-For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
-In general.
-releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
-    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
-For example
-      1. (a)
-        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
-      2. (b)
-        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
-      3. (c)
-        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
-      4. (d)
-        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
-In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
-  21. 5.
-    Open access to data and code
-  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
-  23. Answer: [Yes]
-  24. Justification: [N/A]
-  25. Guidelines:
-    • The answer NA means that paper does not include experiments requiring code.
-    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
-Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
-    • The instructions should contain the exact command and environment needed to run to reproduce the results.
-See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
-    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
-    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
-If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
-    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
-    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
-  26. 6.
-    Experimental setting/details
-  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
-  28. Answer: [Yes]
-  29. Justification: [N/A]
-  30. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
-    • The full details can be provided either with the code, in appendix, or as supplemental material.
-  31. 7.
-    Experiment statistical significance
-  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
-  33. Answer: [Yes]
-  34. Justification: [N/A]
-  35. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
-    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
-    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
-    • The assumptions made should be given (e.g., Normally distributed errors).
-    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
-    • It is OK to report 1-sigma error bars, but one should state it.
-The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
-    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
-    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
-  36. 8.
-    Experiments compute resources
-  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
-  38. Answer: [Yes]
-  39. Justification: [N/A]
-  40. Guidelines:
-    • The answer NA means that the paper does not include experiments.
-    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
-    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
-    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
-  41. 9.
-    Code of ethics
-  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
-  43. Answer: [Yes]
-  44. Justification: [N/A]
-  45. Guidelines:
-    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
-    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
-    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
-  46. 10.
-    Broader impacts
-  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
-  48. Answer: [N/A]
-  49. Justification: [N/A]
-  50. Guidelines:
-    • The answer NA means that there is no societal impact of the work performed.
-    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
-    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
-    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
-However, if there is a direct path to any negative applications, the authors should point it out.
-For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
-On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
-    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
-    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
-  51. 11.
-    Safeguards
-  52.
-Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
-  53. Answer: [N/A]
-  54. Justification: [N/A]
-  55. Guidelines:
-    • The answer NA means that the paper poses no such risks.
-    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
-    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
-    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
-  56. 12.
-    Licenses for existing assets
-  57.
-Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
-  58. Answer: [Yes]
-  59. Justification: [N/A]
-  60. Guidelines:
-    • The answer NA means that the paper does not use existing assets.
-    • The authors should cite the original paper that produced the code package or dataset.
-    • The authors should state which version of the asset is used and, if possible, include a URL.
-    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
-    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
-    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
-For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
-Their licensing guide can help determine the license of a dataset.
-    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
-    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
-  61. 13.
-    New assets
-  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
-  63. Answer: [N/A]
-  64. Justification: [N/A]
-  65. Guidelines:
-    • The answer NA means that the paper does not release new assets.
-    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
-    • The paper should discuss whether and how consent was obtained from people whose asset is used.
-    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
-  66. 14.
-    Crowdsourcing and research with human subjects
-  67.
-Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
-  68. Answer: [N/A]
-  69. Justification: [N/A]
-  70. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
-    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
-  71. 15.
-    Institutional review board (IRB) approvals or equivalent for research with human subjects
-  72.
-Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
-  73. Answer: [N/A]
-  74. Justification: [N/A]
-  75. Guidelines:
-    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
-    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
-If you obtained IRB approval, you should clearly state this in the paper.
-    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
-    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
-  76. 16.
-    Declaration of LLM usage
-  77.
-Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
-Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
-  78. Answer: [N/A]
-  79. Justification: [N/A]
-  80. Guidelines:
-    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
-    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
+Figure 13: Attention Map from post-imputation: Text as query and Vision, Text as key.
+The key tokens are obtained by concatenating tokens from K=2 expert, resulting 100 length of token.
+The darker the color, the lower attention weight it is.
+Figure 14: Attention Map from post-imputation: Vision as query and Audio, Text as key.
+The key tokens are obtained by concatenating tokens from K=2 expert, resulting 100 length of token.
+The darker the color, the lower attention weight it is.
+Figure 15: Attention Map from post-imputation: Audio as query and Vision, Text as key.
+The key tokens are obtained by concatenating tokens from K=2 expert, resulting 100 length of token.
+The darker the color, the lower attention weight it is.
+From Figure 13 14 15, we can clearly observed that the attention map are distinctly different when the key are set to different modalities.
+This indicate that the post-imputation module can dynamically learn, which token are needed depending on characteristics of token and experts.
+For example, the Figure 15 indicates that 25 - 45th tokens of the first expert are more important than others while the text modalities shows that the second experts are more important.
+Similar observation are also shows in 13 14 as well.
+Therefore, we can conclude that the post-imputation can capture expertise of different expert and refine the missing modality accordingly.